### v1

# Be More Diverse than the Most Diverse: Online Selection of Diverse Mixtures of Generative Models

Be More Diverse than the Most Diverse: Online Selection of Diverse Mixtures of Generative Models
Parham Rezaei , Farzan Farnia , Cheuk Ting Li Department of Computer Engineering, Sharif University of Technology, parham.rezaei@sharif.eduDepartment of Computer Science and Engineering, The Chinese University of Hong Kong, farnia@cse.cuhk.edu.hkDepartment of Information Engineering, The Chinese University of Hong Kong, ctli@ie.cuhk.edu.hk
The availability of multiple training algorithms and architectures for generative models requires a selection mechanism to form a single model over a group of well-trained generation models. The selection task is commonly addressed by identifying the model that maximizes an evaluation score based on the diversity and quality of the generated data. However, such a best-model identification approach overlooks the possibility that a mixture of available models can outperform each individual model. In this work, we explore the selection of a mixture of multiple generative models and formulate a quadratic optimization problem to find an optimal mixture model achieving the maximum of kernel-based evaluation scores including kernel inception distance (KID) and Rényi kernel entropy (RKE). To identify the optimal mixture of the models using the fewest possible sample queries, we propose an online learning approach called Mixture Upper Confidence Bound (Mixture-UCB). Specifically, our proposed online learning method can be extended to every convex quadratic function of the mixture weights, for which we prove a concentration bound to enable the application of the UCB approach. We prove a regret bound for the proposed Mixture-UCB algorithm and perform several numerical experiments to show the success of the proposed Mixture-UCB method in finding the optimal mixture of text-based and image-based generative models. The codebase is available at https://github.com/rezaei-parham/mixture-ucb.
1 Introduction
The rapid advancements in generative modeling have created a need for mechanisms to combine multiple well-trained generative models, each developed using different algorithms and architectures, into a single unified model. A common approach for creating such a unified model is to evaluate assessment scores that quantify the diversity and quality of the generated data and then select the model with the highest score. This best-model identification strategy has been widely adopted for the selection of generative models across various domains, including image, text, audio, and video generation.
Figure 1: Visual comparison of the diversity across individual arms and the optimal mixture for images generated using models Kandinsky 3, Stable Diffusion 3-medium, and PixArt-$\alpha$ with the prompt ”Red bird, cartoon style”
Existing model selection frameworks typically perform an offline selection, where they have access to a sufficiently large number of samples from each generative model and estimate the evaluation score based on these samples. However, in many practical scenarios, generating a large sample set from sub-optimal models can be computationally costly, especially if the evaluator can identify their lack of optimality using fewer samples. In such cases, the evaluator can adopt an online learning approach and frame the problem as a multi-armed bandit (MAB) task. An existing approach is successive halving [1, 2, 3],111[2] focused on applying successive halving on hyperparameter optimization for supervised learning, whereas [3] focused on generative models using maximum mean discrepancy [4] as the score. where the models are evaluated using a fixed budget, the worst half of the models are removed, and we repeat the process until one model is left. Also, the recent work by Hu et al. [5] attempts to solve the online model selection problem using an upper confidence bound (UCB) method to identify the generative model with the highest evaluation score. The numerical results in this work indicate the effectiveness of MAB algorithms in reducing sample generation costs from sub-optimal models.
On the other hand, when the model selection task is handled by an online learning algorithm, the algorithm may choose different models at different iterations, resulting in generated data that follow a mixture of the distributions of these generative models. Note that in a standard MAB algorithm, the goal is to eventually converge to a single arm. Successive halving [1, 2, 3] and the standard UCB algorithm adopted by [5] will ultimately select only one generative model after a sufficient number of iterations. However, the diversity scores of the generated data could be higher when the sample generation follows a mixture of models rather than a single model. This observation leads to the following question: could the diversity of generated data be improved by applying MAB algorithms to multiple generative models, if the algorithm aims to find the best mixture of the models?
In this work, we aim to address the above question by finding the optimal mixture $\sum_{i=1}^{m}\alpha_{i}P_{i}$ of $m$ generative models with distributions $P_{1},\ldots,P_{m}$, which would produce a higher evaluation score compared to each individual model. Specifically, we focus on addressing this task in an online learning setting, where we pick a model to generate a sample at each round. To address this problem and develop an MAB algorithm to find the best mixture model, we concentrate on evaluation scores that are quadratic functions of the generated data. As we show in this work, formulating the optimization problem for a quadratic score function results in a quadratic online convex optimization problem that can be efficiently solved using the online gradient descent algorithm. More importantly, we establish a concentration bound for the quadratic function of the mixture weights, which enables us to extend the UCB algorithm for the online selection of the mixture weights.
Specifically, we focus on evaluation scores that reduce to a quadratic function of the generative model’s distribution, including the kernel-based Maximum Mean Discrepancy (MMD) [4], Kernel Inception Distance (KID) [6] and Rényi Kernel Entropy (RKE) [7] scores, as well as the quality-measuring Precision [8, 9] and Density [10] scores, which are linear functions of the generative distribution. Among these scores, RKE provides a reference-free entropy function for assessing the diversity of generated data, making it suitable for quantifying the variety of generated samples. Our mixture-based online learning framework can therefore be applied to find the mixture model with the maximum RKE-based diversity score. Additionally, we can consider a linear combination of RKE with the Precision, Density, or KID quality scores to identify a mixture of models that offers the best trade-off between quality and diversity.
Figure 2: Comparison of our proposed algorithms Mixture-UCB-CAB and Mixture-UCB-OGD with the baseline one arm online algorithms. We plot the diversity-measuring RKE score of the generated data and display 9 (randomly-selected) samples produced in the process of each algorithm.
We perform several numerical experiments to test the application of our proposed Mixture-UCB approach in comparison to the Vanilla-UCB and One-Arm Oracle approaches that tend to generate samples from only one of the available generative models. Our numerical results indicate that the Mixture-UCB algorithms can generate samples with higher RKE diversity scores, and tends to generate samples from a mixture of several generative models when applied to image-based generative models. Also, we test the performance of Mixture-UCB on the KID, Precision, and Density scores, which similarly result in a higher score value for the mixture model found by the Mixture-UCB algorithm. We implement the Mixture-UCB by solving the convex optimization sub-problem at every iteration and also by applying the online gradient descent algorithm at every iteration. In our experiments, both implementations result in satisfactory results and can improve upon learning strategies tending to select only one generative model. The following is a summary of the contributions in this work:
  • Studying the selection task for mixtures of multiple generative models to improve the diversity of generated samples.
  • Proposing an online learning framework to address the mixture selection task for quadratic score functions.
  • Developing the Mixture-UCB-CAB and Mixture-UCB-OGD algorithms to solve the formulated online learning problem.
  • Proving a finite horizon regret bound for Mixture-UCB-CAB which shows the convergence of Mixture-UCB-CAB to the optimal mixture.
  • Designing the Sparse-Mixture-UCB-CAB algorithm for the situation where we want to eventually select a small subset of the models. This is deferred to Appendix B.
  • Presenting numerical results on the improvements in the diversity of generated data by the online selection of a mixture of the generation models.
2 Related Work
Assessment of Generative Models. The evaluation of generative models has been extensively studied, with a focus on both diversity and quality of generated images. Reference-free metrics such as Rényi Kernel Entropy (RKE) [7], VENDI [11], and the FKEA-VENDI [12] measure diversity without relying on ground-truth, while reference-based metrics such as Recall [9] and Coverage [10] assess diversity relative to real data. For image quality evaluation, Density and Precision metrics [10, 9] provide measures based on alignment with a reference distribution. The Wasserstein distance [13] and Fréchet Inception Distance (FID) [14] approximate the distance between real and generated datasets, while Kernel Inception Distance (KID) [6] uses squared maximum mean discrepancy for a kernel-based comparison of distributions. Also, the kernel-based KEN score [15] measures the novelty of generated samples produced by a generative model based on the entropy of their generated sample after canceling the mode-based diversity of a reference distribution.
Multi-Armed Bandit Algorithms. The Multi-Armed Bandit (MAB) problem is a foundational topic in reinforcement learning, where an agent aims to maximize rewards from multiple options (arms) with initially unknown reward distributions [16, 17]. The Upper Confidence Bound (UCB) algorithm [18, 19, 20] is a widely adopted method for addressing the MAB problem, where uncertainty about an arm’s reward is replaced by an optimistic estimate. In generative models, optimism-based bandits have been applied to efficiently identify models with optimal Fréchet Inception Distance (FID) or Inception Score while minimizing data queries [5]. A special case of MAB, the continuum-armed bandit (CAB) problem [21], optimizes a function over continuous inputs, and has been applied to machine learning tasks such as hyperparameter optimization [22, 23]. Recent research explores CABs under more general smoothness conditions like Besov spaces [24], while other works have focused on regret bounds and Lipschitz conditions [25, 26, 27].
Another related reference is informational multi-armed bandits [28], which extends UCB to maximizing the Shannon entropy of a discrete distribution, which is also a metric of diversity. In comparison, the algorithms in this paper can minimize the expectation of any quadratic positive semidefinite function, which not only covers the order-$2$ Rényi entropy for discrete distributions, but also includes the Rényi Kernel Entropy applicable to continuous data. Since the outputs of generative models are generally continuous, [28] is not applicable here. We also note that the recent paper [29] designs a contextual bandit algorithm for the online prompt-based selection of text-guided generative models, where the online learning algorithm adapts the selection of text-based generative models based on their UCB-estimated relevance CLIPScore in response to the current prompt. We note that our work studies a different online learning setting where the goal is to maximize the score of the mixture of unconditional generative models without any input text prompt.
3 Preliminaries
We review several kernel-based performance metrics of generative models.
3.1 Rényi Kernel Entropy
The Rényi Kernel Entropy [7] of the distribution $P$, which measures the diversity of the modes in $P$, is given by $\log(1/\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[k^{2}(X,X^{\prime})])$, where $k$ is a positive definite kernel. [^2] The order-$2$ Rényi entropy for discrete distributions is a special case by taking $k(x,x^{\prime})=\mathbf{1}_{x=x^{\prime}}$. Taking the exponential of the Rényi Kernel Entropy, we have the RKE mode count $1/\mathbb{E}[k^{2}(X,X^{\prime})])$ [7], which is an estimate of the number of modes. Maximizing the RKE mode count is equivalent to minimizing the following loss
$\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[k^{2}(X,X^{\prime})].$ (1)
3.2 Maximum Mean Discrepancy and Kernel Inception Distance
The (squared) maximum mean discrepancy (MMD) [4] between distributions $P,Q$, which measures the distance between $P$ and $Q$, can be written as
$\mathbb{E}[k(X,X^{\prime})]+\mathbb{E}[k(Y,Y^{\prime})]-2\mathbb{E}[k(X,Y)],$ (2)
where $X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P$ and $Y,Y^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}Q$, and $k$ is a positive definite kernel. Suppose $P$ is the distribution of samples from a generative model, and $Q$ is a reference distribution. Minimizing the MMD can ensure that $P$ is close to $Q$. The Kernel Inception Distance (KID) [6], a popular quality metric for image generative models, is obtained by first passing $P$ and $Q$ through the Inception network [30], and then computing their MMD, i.e., we have
$\mathbb{E}[k(\psi(X),\psi(X^{\prime}))]+\mathbb{E}[k(\psi(Y),\psi(Y^{\prime}))]-2\mathbb{E}[k(\psi(X),\psi(Y))],$ (3)
where $\psi$ is the mapping from $x$ to its Inception representation.
4 Optimal Mixtures of Generative Models
RKE (1), MMD (2) and KID (3) can all be written as a loss function in the following form
$L(P):=\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[\kappa(X,X^{\prime})]+\mathbb{E}_{X\sim P}[f(X)],$ (4)
where $\kappa:\mathcal{X}^{2}\to\mathbb{R}$ is a positive semidefinite kernel, and $f:\mathcal{X}\to\mathbb{R}$ is a function. For (1), we take $\kappa(x,x^{\prime})=k^{2}(x,x^{\prime})$ (the square of a kernel is still a kernel) and $f(x)=0$. For (2), we take $\kappa(x,x^{\prime})=k(x,x^{\prime})$ and $f(x)=-2\mathbb{E}_{Y\sim Q}[k(x,Y)]$ (the constant term $\mathbb{E}[k(Y,Y^{\prime})]$ does not matter). For KID (3), we take $\kappa(x,x^{\prime})=k(\psi(x),\psi(x^{\prime}))$ and $f(x)=-2\mathbb{E}_{Y\sim Q}[k(\psi(x),\psi(Y))]$. Note that any convex combinations of (1), (2) and (3) is still in the form (4).
Suppose we are given $m$ generative models, where model $i$ generates samples from the distribution $P_{i}$. If our goal is merely to find the model that minimize the loss (4), we should select $\mathrm{argmin}_{i}L(P_{i})$. Nevertheless, for diversity metrics such as RKE, it is possible that a mixture of the models will give a better diversity. Assume that the mixture weight of model $i$ is $\alpha_{i}\in[0,1]$, where $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{m})$ is a probability vector. The loss of the mixture distribution $\sum_{i=1}^{m}\alpha_{i}P_{i}$ can then be expressed as
$L(\boldsymbol{\alpha})$ $:=L\Big{(}\sum_{i=1}^{m}\alpha_{i}P_{i}\Big{)}=\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}+\mathbf{f}^{\top}\boldsymbol{\alpha},$
$\mathbf{K}:=(\mathbb{E}_{X\sim P_{i},\,X^{\prime}\sim P_{j}}[\kappa(X,X^{\prime})])_{i,j\in[m]}\in\mathbb{R}^{m\times m},\quad\mathbf{f}:=(\mathbb{E}_{X\sim P_{i}}[f(X)])_{i=1}^{m}\in\mathbb{R}^{m}.$
Given $\mathbf{K},\mathbf{f}$, the probability vector $\boldsymbol{\alpha}$ minimizing $L(\boldsymbol{\alpha})$ can be found via a convex quadratic program.
In practice, we do not know the precise $\mathbf{K},\mathbf{f}$, and have to estimate them using samples. Suppose we have the samples $x_{i,1},\ldots,x_{i,n_{i}}$ from the distribution $P_{i}$ for $i=1,\ldots,m$, where $n_{i}$ is the number of observed samples from model $i$. Write $\mathbf{x}:=(x_{i,a})_{i\in[m],\,a\in[n_{i}]}$. We approximate the true mixture distribution $\sum_{i=1}^{m}\alpha_{i}P_{i}$ by the empirical mixture distribution $\sum_{i=1}^{m}\frac{\alpha_{i}}{n_{i}}\sum_{a=1}^{n_{i}}\delta_{x_{i,a}}$, where we assign a weight $\alpha_{i}/n_{i}$ to samples $x_{i,a}$ from model $i$, and $\delta_{x_{i,a}}$ denotes the degenerate distribution at $x_{i,a}$. We then approximate $L(\boldsymbol{\alpha})$ by the sample loss
$\hat{L}(\boldsymbol{\alpha};\mathbf{x})$ $:=L\Big{(}\sum_{i=1}^{m}\frac{\alpha_{i}}{n_{i}}\sum_{a=1}^{n_{i}}\delta_{x_{i,a}}\Big{)}=\boldsymbol{\alpha}^{\top}\hat{\mathbf{K}}(\mathbf{x})\boldsymbol{\alpha}+\hat{\mathbf{f}}(\mathbf{x})^{\top}\boldsymbol{\alpha},$ (5)
$\hat{\mathbf{K}}(\mathbf{x}):=\Big{(}\frac{1}{n_{i}n_{j}}\sum_{a=1}^{n_{i}}\sum_{b=1}^{n_{j}}\kappa(x_{i,a},x_{j,b})\Big{)}_{i,j}\in\mathbb{R}^{m\times m},\quad\hat{\mathbf{f}}(\mathbf{x}):=\Big{(}\frac{1}{n_{i}}\sum_{a=1}^{n_{i}}f(x_{i,a})\Big{)}_{i=1}^{m}\in\mathbb{R}^{m}.$
The minimization of $\hat{L}(\boldsymbol{\alpha};\mathbf{x})$ over probability vectors $\boldsymbol{\alpha}$ is still a convex quadratic program.
5 Online Selection of Optimal Mixtures – Mixture Multi-Armed Bandit
Suppose we are given $m$ generative models, but we do not have any prior information about them. Our goal is to use these models to generate a collection of samples $(x^{(t)})_{i\in[T]}$ in $T$ rounds that minimizes the loss (4) $L(\hat{P}^{(T)})$ at the empirical distribution $\hat{P}^{(T)}=T^{-1}\sum_{t=1}^{T}\delta_{x^{(t)}}$. We have
$L(\hat{P}^{(T)})=\frac{1}{T^{2}}\sum_{s,t\in[T]}\kappa(x^{(s)},x^{(t)})+\frac{1}{T}\sum_{t=1}^{T}f(x^{(t)}).$
If we are told by an oracle the optimal mixture $\boldsymbol{\alpha}^{*}$ that minimizes the loss $L(\boldsymbol{\alpha})$, then we should generate samples according to this mixture distribution, giving $\approx\alpha^{*}_{i}T$ samples from model $i$. We call this the mixture oracle scenario. Nevertheless, in reality, we do not know $\mathbf{K},\mathbf{f}$, and cannot compute $\boldsymbol{\alpha}^{*}$ exactly. Instead, we have to approximate $\boldsymbol{\alpha}^{*}$ by minimizing the sample loss $\hat{L}(\boldsymbol{\alpha};\mathbf{x})$ (5). However, we do not have the samples $\mathbf{x}$ at the beginning in order to compute $\hat{L}(\boldsymbol{\alpha};\mathbf{x})$, so we have to generate some samples first. Yet, to generate these initial samples, we need an estimate of $\boldsymbol{\alpha}^{*}$, or else those samples may have a suboptimal empirical distribution and affect our final loss $L(\hat{P}^{(T)})$, or we will have to discard those initial samples which results in wastage.
This “chicken and egg” problem is naturally solved by an online learning approach via multi-armed bandit. At time $t=1,\ldots,T$, we choose and pull an arm $b^{(t)}\in[m]$ (i.e., generate a sample from model $b^{(t)}$), and obtain a sample $x^{(t)}$ from the distribution $P_{b^{(t)}}$. The choice $b^{(t)}$ can depend on all previous samples $x^{(1)},\ldots,x^{(t-1)}$. Unlike conventional multi-armed bandit where the goal is to maximize the total reward over $T$ rounds, here we minimize the loss $L(\hat{P}^{(T)})$ which involve cross terms $\kappa(x^{(s)},x^{(t)})$ between samples at different rounds. Note that if $\kappa(x,x^{\prime})=0$, then this reduces to the conventional multi-armed bandit setting by taking $f(x)$ to be the negative reward of the sample $x$. In the following subsections, we will propose two new algorithms that are generalizations of the upper confidence bound (UCB) algorithm for multi-armed bandit [18, 19].
5.1 Mixture Upper Confidence Bound – Continuum-Armed Bandit
Let $n_{i}^{(t)}$ be the number of times arm $i$ has been pulled up to time $t$. Let $\mathbf{x}^{(t)}:=(x_{i,a})_{i\in[m],\,a\in[n_{i}^{(t)}]}$ be the observed samples up to time $t$. We focus on bounded loss functions, and assume that $\kappa:\mathcal{X}^{2}\to[\kappa_{0},\kappa_{1}]$ and $f:\mathcal{X}\to[f_{0},f_{1}]$ are bounded. Let $\Delta_{\kappa}:=\kappa_{1}-\kappa_{0}$, $\Delta_{f}:=f_{1}-f_{0}$. Define the sensitivity of $L$ as $\Delta_{L}:=2\Delta_{\kappa}+\Delta_{f}$.
We now present the mixture upper confidence bound – continuum-armed bandit (Mixture-UCB-CAB) algorithm. It has a parameter $\beta>1$. It treats the online selection problem as multi-armed bandit with infinitely many arms similar to the continuum-armed bandit settings in [21, 31]. Each arm is a probability vector $\boldsymbol{\alpha}$. By pulling the arm $\boldsymbol{\alpha}$, we generate a sample from a randomly chosen model, where model $i$ is chosen with probability $\alpha_{i}$. Refer to Algorithm 1.
Algorithm 1 Mixture-UCB-CAB
1:Input: $m$ generative arms, number of rounds $T$
2:Output: Gathered samples $\mathbf{x}^{(T)}$
3:for $t\in\{0,\ldots,m-1\}$  do
4:     Pull arm $t+1$ at time $t+1$ to obtain sample $x_{t+1,1}\sim P_{t+1}$. Set $n_{t+1}^{(m)}=1$.
5:end for
6:for $t\in\{m,\ldots,T-1\}$  do
7:     Compute an estimate of the optimal mixture distribution via the convex quadratic program:
$\boldsymbol{\alpha}^{(t)}:=\mathrm{argmin}_{\boldsymbol{\alpha}}\big{(}\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}\big{)},$ (6)
where the minimization is over probability vectors $\boldsymbol{\alpha}$, and $\boldsymbol{\epsilon}^{(t)}\in\mathbb{R}^{m}$ is defined as
$\epsilon_{i}^{(t)}:=\Delta_{L}\sqrt{(\beta\log t)/(2n_{i}^{(t)})}+\Delta_{\kappa}/n_{i}^{(t)}.$ (7)
8:     Generate the arm index $b^{(t+1)}\in[m]$ at random with $\mathbb{P}(b^{(t+1)}=i)=\alpha_{i}^{(t)}$.
9:     Pull arm $b=b^{(t+1)}$ at time $t+1$ to obtain a new sample $x_{b,n_{b}^{(t)}+1}\sim P_{b}$. Set $n_{b}^{(t+1)}=n_{b}^{(t)}+1$ and $n_{j}^{(t+1)}=n_{j}^{(t)}$ for $j\neq b$33footnotemark: 3
10:end forreturn samples $\mathbf{x}^{(T)}$
Similar to UCB, Mixture-UCB-CAB finds a lower confidence bound $\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}$ of the true loss $L(\boldsymbol{\alpha})$ at each round. To justify the expressions (6), (7), we prove that $\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}$ in (6) lower-bounds $L(\boldsymbol{\alpha})$ with probability at least $1-t^{-\beta}$. The proof is given in Appendix A.1.
Theorem 1.
Fix a probability vector $\boldsymbol{\alpha}$ [^4] Theorem 1 holds for a fixed $\boldsymbol{\alpha}$. A worst-case bound that simultaneously holds for every $\boldsymbol{\alpha}$ is in Lemma 1. Suppose we have samples $x_{i,1},\ldots,x_{i,n_{i}}$ from the distribution $P_{i}$ for $i=1,\ldots,m$, where $n_{i}$ is the number of observed samples from model $i$. For $\delta>0$,
$\mathbb{P}\big{(}\hat{L}(\boldsymbol{\alpha};\mathbf{x})-L(\boldsymbol{\alpha})\geq\boldsymbol{\epsilon}(\delta)^{\top}\boldsymbol{\alpha}\big{)}\leq\delta,\quad\mathbb{P}\big{(}L(\boldsymbol{\alpha})-\hat{L}(\boldsymbol{\alpha};\mathbf{x})\geq\boldsymbol{\epsilon}(\delta)^{\top}\boldsymbol{\alpha}\big{)}\leq\delta,$
where $\boldsymbol{\epsilon}(\delta):=\big{(}\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\big{)}_{i\in[m]}$.
We now prove that Mixture-UCB-CAB gives an expected loss $\mathbb{E}[L(\hat{P}^{(T)})]$ that converges to the optimal loss $\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha})$ by bounding their gap. This means that Mixture-UCB-CAB is a zero-regret strategy by treating $\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha})$ as the average regret per round. [^5] To justify calling $R:=\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha})$ the average regret per round, note that when $\kappa(x,x^{\prime})=0$ and $f(x)=-r(x)$ where $r(x)$ is the reward of the sample $x$, i.e., the loss $L(P)=\mathbb{E}_{X\sim P}[f(X)]$ is linear, $T(\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha}))=T\max_{i\in[m]}\mathbb{E}_{X\sim P_{i}}[r(X)]-\mathbb{E}[\sum_{t=1}^{T}r(x^{(t)})]$ indeed reduces to the conventional notion of regret. So $R$ can be regarded as the quadratic generalization of regret. The proof is given in Appendix A.2.
Theorem 2.
Suppose $m\geq 2$, $\beta\geq 4$. Consider bounded quadratic loss function (4) with $\kappa$ being positive semidefinite. Let $\hat{P}^{(T)}$ be the empirical distribution of the first $T\geq 2$ samples $\mathbf{x}^{(T)}$ given by Mixture-UCB-CAB. Then the gap between the expected loss and the optimal loss is bounded by
$\mathbb{E}\big{[}L(\hat{P}^{(T)})\big{]}-\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha})$ $\leq 4\Delta_{L}\sqrt{\frac{\beta m\log T}{T}}.$
When $\kappa(x,x^{\prime})=0$, Mixture-UCB-CAB reduces to the conventional UCB, and Theorem 2 coincides with the $O(\sqrt{(m\log T)/T})$ distribution-free bound on the regret per round of conventional UCB [20]. Since there is a $\Omega(\sqrt{m/T})$ minimax lower bound on the regret per round even for conventional multi-armed bandit without the quadratic kernel term [20, Theorem 3.4], Theorem 2 is tight up to a logarithmic factor.
The main difference between Mixture-UCB-CAB and conventional UCB is that we choose a mixture of arms in (6) given by the probability vector $\boldsymbol{\alpha}$, instead of a single arm. A more straightforward application of UCB would be to simply find the single arm that minimizes the lower bound in (6), i.e., we restrict $\boldsymbol{\alpha}=\mathbf{e}_{i}$ for some $i\in[m]$, where $\mathbf{e}_{i}$ is the $i$-th basis vector, and minimize (6) over $i$ instead. We call this Vanilla-UCB. Vanilla-UCB fails to take into account the possibility that a mixture may give a smaller loss than every single arm. In the long run, Vanilla-UCB converges to pulling the best single arm instead of the optimal mixture. Vanilla-UCB will be used as a baseline to be compared with Mixture-UCB-CAB, and another new algorithm presented in the next section.
Mixture-UCB-CAB can be extended to the Sparse-Mixture-UCB-CAB algorithm which eventually select only a small number of models. This can be useful if there is a subscription cost for each model. Refer to Appendix B for discussions.
55footnotetext: We may also consider the scenario where each pull gives a batch of $l$ samples instead of only one sample. In this case, we will have $x_{b,n_{b}^{(t-1)}+1},\ldots,x_{b,n_{b}^{(t-1)}+l}\sim P_{b}$ and $n_{b}^{(t)}=n_{b}^{(t-1)}+l$.
5.2 Mixture Upper Confidence Bound – Online Gradient Descent
We present an alternative to Mixture-UCB-CAB, called the mixture upper confidence bound – online gradient descent (Mixture-UCB-OGD) algorithm, inspired by the online gradient descent algorithm [32]. It also has a parameter $\beta>1$. Refer to Algorithm 2.
Algorithm 2 Mixture-UCB-OGD
1:Input: $m$ generative arms, number of rounds $T$
2:Output: Gathered samples $\mathbf{x}^{(T)}$
3:for $t\in\{0,\ldots,m-1\}$  do
4:     Pull arm $t+1$ at time $t+1$ to obtain sample $x_{t+1,1}\sim P_{t+1}$
5:end for
6:for $t\in\{m,\ldots,T-1\}$  do
7:     Compute the gradient
$\mathbf{h}^{(t)}$ $:=\nabla_{\boldsymbol{\alpha}}\left(\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}\right)\Big{|}_{\boldsymbol{\alpha}=\mathbf{n}^{(t)}/t}\;=\;\frac{2}{t}\hat{\mathbf{K}}(\mathbf{x}^{(t)})\mathbf{n}^{(t)}+\hat{\mathbf{f}}(\mathbf{x}^{(t)})-\boldsymbol{\epsilon}^{(t)},$ (8)
where $\mathbf{n}^{(t)}:=(n_{i}^{(t)})_{i\in[m]}\in\mathbb{R}^{m}$, and $\boldsymbol{\epsilon}^{(t)}$ is defined as in Mixture-UCB-CAB
8:     Pull arm $b=b^{(t+1)}:=\mathrm{argmin}_{i}h_{i}^{(t)}$ at time $t+1$ to obtain a new sample $x_{b,n_{b}^{(t)}+1}\sim P_{b}$.
9:end forreturn samples $\mathbf{x}^{(T)}$
Mixture-UCB-CAB and Mixture-UCB-OGD can both be regarded as generalizations of the original UCB algorithm, in the sense that they reduce to UCB when $\kappa(x,x^{\prime})=0$. If we remove the $\frac{2}{t}\hat{\mathbf{K}}(\mathbf{x})\mathbf{n}^{(t)}$ term in (8), then Mixture-UCB-OGD becomes the same as UCB.
Both Mixture-UCB-CAB and Mixture-UCB-OGD attempt to make the “proportion vector” $\mathbf{n}^{(t)}/t$ (note that $n_{i}^{(t)}/t$ is the proportion of samples from model $i$) approach the optimal mixture $\boldsymbol{\alpha}^{*}$ that minimizes $L(\boldsymbol{\alpha})$, but they do so in different manners. Mixture-UCB-CAB first computes the estimate $\boldsymbol{\alpha}^{(t)}$ after time $t$, then approaches $\boldsymbol{\alpha}^{(t)}$ by pulling an arm randomly chosen from the distribution $\boldsymbol{\alpha}^{(t)}$. Mixture-UCB-OGD estimates the gradient $\mathbf{h}^{(t)}$ of the loss function at the current proportion vector $\mathbf{n}^{(t)}/t$, and pulls an arm that results in the steepest descent of the loss.
An advantage of Mixture-UCB-OGD is that the computation of gradient (8) is significantly faster than the quadratic program (6) in Mixture-UCB-CAB. The running time complexity of Mixture-UCB-OGD is $O(T^{2}+Tm^{2})$. [^6] To update $\hat{\mathbf{K}}(\mathbf{x}^{(t)})$ after a new sample $x^{\prime}$ is obtained, we only need to compute $\kappa(x,x^{\prime})$ for each existing sample $x$, and add their contributions to the corresponding entries in $\hat{\mathbf{K}}(\mathbf{x}^{(t)})$, requiring a computational time that is linear with the number of existing samples. Nevertheless, a regret bound for Mixture-UCB-OGD similar to Theorem 2 seems to be difficult to derive, and is left for future research.
6 Numerical Results
We experiment on various scenarios to showcase the performance of our proposed algorithms. The experiments involve the following algorithms:
  • Mixture Oracle. In the mixture oracle algorithm (Section 5), an oracle tells us the optimal mixture $\boldsymbol{\alpha}^{*}$ in advance, and we pull arms randomly according to this distribution. The optimal mixture is calculated by solving the quadratic optimization in Section 4 on a large number of samples for each arm. The number of chosen samples varies based on the experiments. This is an unrealistic setting that only serves as a theoretical upper bound of the performance of any online algorithm. A realistic algorithm that performs close to the mixture oracle would be almost optimal.
  • One-Arm Oracle. An oracle tells us the optimal single arm in advance, and we keep pulling this arm. This is an unrealistic setting. If our algorithms outperform the one-arm oracle, this will show that the advantage of pulling a mixture of arms (instead of a single arm) can be realistically achieved via online algorithms.
  • Vanilla-UCB. A direct application of UCB mentioned near the end of Section 5.1. This serves as a baseline for the purpose of comparison.
  • Successive Halving. The Success Halving algorithm [1, 2, 3] which serves as a second baseline for comparison.
  • Mixture-UCB-CAB. The mixture upper confidence bound – continuum-armed bandit algorithm proposed in Section 5.1.
  • Mixture-UCB-OGD. The mixture upper confidence bound – online gradient descent algorithm proposed in Section 5.2.
Experiments Setup.
We use DINOv2-ViT-L/14 [33] for image feature extraction, as recommended by [34], and use RoBERTa [35] as text encoder. Detailed explanation of the setup for each experiment is presented in Section C.1.
6.1 Optimal Mixture for Diversity and Quality via KID
We conduct three experiments to evaluate our method using the Kernel Inception Distance (KID) metric. In the first experiment, we use five distinct generative models: LDM [36], StyleGAN-XL [37], Efficient-vdVAE [38], InsGen [39], and StyleNAT [40], all trained on the FFHQ dataset [41]. In the second experiment, we use generated images from four models [^7] FFHQ and LSUN-Bedroom datasets were downloaded from the dgm-eval repository [34] (licensed under MIT license): https://github.com/layer6ai-labs/dgm-eval.: StyleGAN [41], Projected GAN [42], iDDPM [43], and Unleashing Transformers [44], all trained on the LSUN-Bedroom dataset [45]. This experiment follows a similar setup to the first. In the final experiment, we employ the Truncation Method [46, 41] to generate diversity-controlled images centered on eight randomly selected points, using StyleGAN2-ADA [47], also trained on the FFHQ dataset. Figure 3 demonstrates that the mixture of generators achieves a better KID score compared to individual models. Additionally, the two Mixture-UCB algorithms consistently outperform the baselines.
(a) FFHQ generators
(b) LSUN-Bedroom generators
(c) FFHQ truncated generators
Figure 3: Performance comparison of online algorithms for the KID metric across FFHQ, LSUN-Bedroom, and FFHQ Truncated generators.
6.2 Optimal Mixture for Diversity via RKE
We use the RKE Mode Count [7] as an evaluation metric to show the effect of mixing the models on the diversity and the advantage of our algorithms Mixture-UCB-CAB and Mixture-UCB-OGD. The score in the plots is the RKE Mode Count, written as RKE for brevity.
(a) FFHQ truncated generators
(b) AFHQ truncated generators
Figure 4: Performance comparison of online algorithms based on the RKE metric for Simulator Unconditional Generative Models.
Synthetic Unconditional Generative Models
We conduct two experiments on diversity-limited generative models. First, we use eight center points with a truncation value of 0.3 to generate images using StyleGAN2-ADA, trained on the FFHQ dataset. In the second experiment, we apply the same model, trained on the AFHQ Cat dataset [48], with a truncation value of 0.4. As shown in Figure 4, the mixture achieves a higher RKE score, and our algorithms consistently get a higher RKE value. The increase in diversity is visually depicted in Figures 8 and 9.
Text to Image Generative Models
Figure 5: Visual comparison of the diversity across individual arms and the optimal mixture for Dog Breed Generators and Style-Specific Generators.
Given that many text-to-image generative models are trained on different datasets, we expect a more pronounced increase in diversity when using a mixture of these models, compared to unconditional generative models trained on the same dataset, such as FFHQ.
We explore three scenarios to evaluate the performance of our algorithms. First, we simulate when models struggle to capture different styles. To address this, we use Stable Diffusion XL [49] with specific prompts to create three car image generators with distinct styles: realistic, surreal, and cartoon. In the second experiment, recognizing the importance of diversity in generative models for design tasks, we use four models—FLUX.1-Schnell [50], Kandinsky 3.0 [51], PixArt-$\alpha$ [52], and Stable Diffusion XL—to generate images of the object “Sofa”. In a similar manner, we generate red cartoonish bird images using Kandinsky 3.0, Stable Diffusion 3 [53], and PixArt-$\alpha$. The increase in diversity is shown in Figure 1. The performance of our online algorithms and the comparison of samples generated by them with the baselines are shown in Figure 2. Finally, in the last experiment, we use Stable Diffusion XL to simulate models generating images of different dog breeds: Bulldog, German Shepherd, and Poodle, respectively. This illustrates the challenge of generating diverse object types with text-to-image models. Figure 5 illustrates the impact of using a mixture of models in the first and last experiments. The improvement in diversity is evident both visually and quantitatively, as reflected in the RKE Scores. As shown in Figure 6, our online algorithms consistently outperform others in generating more diverse samples.
(a) Dog breed generators
(b) Sofa image generators
(c) Style-specific generators
Figure 6: Performance comparison of online algorithms using the RKE metric for text-to-image generative models.
Text Generative Models
We utilize the OpenLLMText dataset [54], which comprises 60,000 human texts rephrased paragraph by paragraph using the GPT2-XL [55], LLaMA-7B [56], and PaLM [57] models. To extract textual features, we employ the RoBERTa Text Encoder. As shown in Figure 11(a) in Section C.3, the results demonstrate the advantage of our online algorithms, suggesting that our method applies not only to image generators, but also to text generators.
6.3 Optimal Mixture for Diversity and Quality via RKE and Precision/Density
Using RKE, we focus solely on the diversity of the models without accounting for their quality. To address this, we apply our methodology to both RKE and Precision [9], as well as RKE and Density [10]. We conduct experiments in which quality is a key consideration. We use four models: three are StyleGAN2-ADA models trained on the FFHQ dataset, each generating images with a truncation of 0.3 around randomly selected center points. The fourth model is StyleGAN2-ADA trained on CIFAR-10 [58]. The FFHQ dataset is used as the reference dataset, therefore, the model trained on CIFAR-10 acts as a low quality generator. Figures 7 and 12 demonstrate the ability of our algorithms to find optimal mixtures with a higher diversity/quality score.
(a) RKE w/Precision
(b) RKE w/Density
Figure 7: Performance comparison of online algorithms using the combination of RKE with Precision and RKE with Density metrics.
7 Conclusion
We studied the online selection from several generative models, where the online learner aims to generate samples with the best overall quality and diversity. While standard multi-armed bandit (MAB) algorithms aim to converge to one arm and select one generative model, we highlighted the fact that a mixture of generative models could achieve a higher score compared to each individual model. We proposed the Mixture-UCB-CAB and Mixture-UCB-OGD online learning algorithms to find the optimal mixture. Our experiments suggest the usefulness of the algorithm in improving the performance scores over individual arms. Extending the algorithm to conditional and text-based generative models is a topic for future exploration. In addition, the application of the algorithm to other data domains, including text, audio, and video, is an interesting future direction.
Appendix A Proofs
A.1 Proof of Theorem 1
Consider any $\tilde{x}_{i,a}\in\mathcal{X}$. Let $\tilde{\mathbf{x}}$ be the samples which are identical to $\mathbf{x}$ except that one entry $x_{i,a}$ is changed to $\tilde{x}_{i,a}$. We have
$\left|\hat{L}(\boldsymbol{\alpha};\tilde{\mathbf{x}})-\hat{L}(\boldsymbol{\alpha};\mathbf{x})\right|$
$=\bigg{|}\frac{\alpha_{i}}{n_{i}}\left(f(\tilde{x}_{i,a})-f(x_{i,a})\right)+2\sum_{(j,b)\neq(i,a)}\frac{\alpha_{i}\alpha_{j}}{n_{i}n_{j}}\left(\kappa(\tilde{x}_{i,a},x_{j,b})-\kappa(x_{i,a},x_{j,b})\right)$
$\;\;\;\;\;+\frac{\alpha_{i}^{2}}{n_{i}^{2}}\left(\kappa(\tilde{x}_{i,a},\tilde{x}_{i,a})-\kappa(x_{i,a},x_{i,a})\right)\bigg{|}$
$\leq\frac{\alpha_{i}}{n_{i}}\Delta_{f}+2\sum_{(j,b)\neq(i,a)}\frac{\alpha_{i}\alpha_{j}}{n_{i}n_{j}}\Delta_{\kappa}+\frac{\alpha_{i}^{2}}{n_{i}^{2}}\Delta_{\kappa}$
$\leq\frac{\alpha_{i}}{n_{i}}(\Delta_{f}+2\Delta_{\kappa})$
$=\frac{\alpha_{i}}{n_{i}}\Delta_{L}.$
By McDiarmid’s inequality,
$\mathbb{P}\left(\hat{L}(\boldsymbol{\alpha};\mathbf{x})-\mathbb{E}[\hat{L}(\boldsymbol{\alpha};\mathbf{x})]\geq\epsilon\right)$ $\leq\exp\left(-\frac{2\epsilon^{2}}{\sum_{i=1}^{m}\sum_{a=1}^{n_{i}}(\frac{\alpha_{i}}{n_{i}}\Delta_{L})^{2}}\right)$
$=\exp\left(-\frac{2\epsilon^{2}}{\Delta_{L}^{2}\sum_{i=1}^{m}\alpha_{i}^{2}/n_{i}}\right).$
Note that
$\left|L(\boldsymbol{\alpha})-\mathbb{E}[\hat{L}(\boldsymbol{\alpha};\mathbf{x})]\right|$
$=\bigg{|}\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[\kappa(X,X^{\prime})]-\mathbb{E}\bigg{[}\sum_{(i,j)\in[m]^{2}}\frac{\alpha_{i}\alpha_{j}}{n_{i}n_{j}}\sum_{a=1}^{n_{i}}\sum_{b=1}^{n_{j}}\kappa(x_{i,a},x_{j,b})\bigg{]}\bigg{|}$
$=\bigg{|}\sum_{i=1}^{m}\frac{\alpha_{i}^{2}}{n_{i}^{2}}\sum_{a=1}^{n_{i}}\left(\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[\kappa(X,X^{\prime})]-\mathbb{E}[\kappa(x_{i,a},x_{i,a})]\right)\bigg{|}$
$\leq\sum_{i=1}^{m}\frac{\alpha_{i}^{2}}{n_{i}}\Delta_{\kappa}.$
Hence, for $\delta>0$,
$\mathbb{P}\left(\hat{L}(\boldsymbol{\alpha};\mathbf{x})-L(\boldsymbol{\alpha})\geq\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2}\sum_{i=1}^{m}\frac{\alpha_{i}^{2}}{n_{i}}}+\Delta_{\kappa}\sum_{i=1}^{m}\frac{\alpha_{i}^{2}}{n_{i}}\right)\leq\delta.$
The result follows from
$\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2}\sum_{i=1}^{m}\frac{\alpha_{i}^{2}}{n_{i}}}+\Delta_{\kappa}\sum_{i=1}^{m}\frac{\alpha_{i}^{2}}{n_{i}}$
$\leq\Delta_{L}\sum_{i=1}^{m}\sqrt{\frac{\log(1/\delta)}{2}\frac{\alpha_{i}^{2}}{n_{i}}}+\Delta_{\kappa}\sum_{i=1}^{m}\frac{\alpha_{i}}{n_{i}}$
$=\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\right)\alpha_{i}.$
The other direction of the inequality is similar.
A.2 Proof of Theorem 2
Before we prove Theorem 2, we first prove a worst case concentration bound.
Lemma 1.
Let $T\geq 2$, and $x_{i,1},x_{i,2},\ldots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P_{i}$ for $i\in[m]$. Let $\Delta_{L}:=2\Delta_{\kappa}+\Delta_{f}$. For $n_{1},\ldots,n_{m}\in[T]$, let $\mathbf{x}^{(n_{i})_{i}}:=(x_{i,a})_{i\in[m],\,a\in[n_{i}]}$. Fix any $\delta>0$. With probability at least $1-\delta$, we have
$L(\boldsymbol{\alpha})-\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(n_{i})_{i}})$
$\leq\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{1}{2n_{i}}\log\frac{m^{2}T^{2}}{2\delta}}+\frac{\Delta_{\kappa}}{n_{i}}\right)\alpha_{i}.$
for every $n_{1},\ldots,n_{m}\in[T]$ and probability vector $\boldsymbol{\alpha}$. The same holds for $\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(n_{i})_{i}})-L(\boldsymbol{\alpha})$ instead of $L(\boldsymbol{\alpha})-\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(n_{i})_{i}})$.
Proof: Fix any $n_{1},\ldots,n_{m}\in[T]$, and write $\mathbf{x}=\mathbf{x}^{(n_{i})_{i}}$. We have
$L(\boldsymbol{\alpha})-\hat{L}(\boldsymbol{\alpha};\mathbf{x})$
$=\sum_{i,j}\alpha_{i}\alpha_{j}\left(\frac{\mathbf{f}_{i}+\mathbf{f}_{j}}{2}+\mathbf{K}_{i,j}-\frac{\hat{\mathbf{f}}_{i}(\mathbf{x})+\hat{\mathbf{f}}_{j}(\mathbf{x})}{2}-\hat{\mathbf{K}}_{i,j}(\mathbf{x})\right).$
For $i=j$, applying Theorem 1 on $\boldsymbol{\alpha}$ being the $i$-th basis vector,
$\mathbb{P}\left(\mathbf{f}_{i}+\mathbf{K}_{i,i}-\hat{\mathbf{f}}_{i}(\mathbf{x})-\hat{\mathbf{K}}_{i,i}(\mathbf{x})\geq\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\right)\leq\delta.$ (9)
For $i\neq j$, we will use similar arguments as Theorem 1. Let $\tilde{\mathbf{x}}$ be the samples which are identical to $\mathbf{x}$ except that one entry $x_{i,a}$ of the $i$-th arm is changed to $\tilde{x}_{i,a}$. We have
$\left|\frac{\hat{\mathbf{f}}_{i}(\tilde{\mathbf{x}})+\hat{\mathbf{f}}_{j}(\tilde{\mathbf{x}})}{2}+\hat{\mathbf{K}}_{i,j}(\tilde{\mathbf{x}})-\frac{\hat{\mathbf{f}}_{i}(\mathbf{x})+\hat{\mathbf{f}}_{j}(\mathbf{x})}{2}-\hat{\mathbf{K}}_{i,j}(\mathbf{x})\right|$
$=\bigg{|}\frac{1}{2n_{i}}\left(f(\tilde{x}_{i,a})-f(x_{i,a})\right)+\frac{1}{n_{i}n_{j}}\sum_{b=1}^{n_{j}}\left(\kappa(\tilde{x}_{i,a},x_{j,b})-\kappa(x_{i,a},x_{j,b})\right)\bigg{|}$
$\leq\frac{\Delta_{f}}{2n_{i}}+\frac{\Delta_{\kappa}}{n_{i}}$
$=\frac{\Delta_{L}}{2n_{i}}.$
Note that $\mathbb{E}[\frac{\hat{\mathbf{f}}_{i}(\mathbf{x})+\hat{\mathbf{f}}_{j}(\mathbf{x})}{2}+\hat{\mathbf{K}}_{i,j}(\mathbf{x})]=\frac{\mathbf{f}_{i}+\mathbf{f}_{j}}{2}-\mathbf{K}_{i,j}$. By McDiarmid’s inequality,
$\mathbb{P}\left(\frac{\mathbf{f}_{i}+\mathbf{f}_{j}}{2}+\mathbf{K}_{i,j}-\frac{\hat{\mathbf{f}}_{i}(\mathbf{x})+\hat{\mathbf{f}}_{j}(\mathbf{x})}{2}-\hat{\mathbf{K}}_{i,j}(\mathbf{x})\geq\epsilon\right)$
$\leq\exp\left(-\frac{2\epsilon^{2}}{n_{i}\left(\frac{\Delta_{L}}{2n_{i}}\right)^{2}+n_{j}\left(\frac{\Delta_{L}}{2n_{j}}\right)^{2}}\right)$
$=\exp\left(-\frac{8\epsilon^{2}}{\Delta_{L}^{2}\left(n_{i}^{-1}+n_{j}^{-1}\right)}\right).$
Hence,
$\mathbb{P}\left(\frac{\mathbf{f}_{i}+\mathbf{f}_{j}}{2}+\mathbf{K}_{i,j}-\frac{\hat{\mathbf{f}}_{i}(\mathbf{x})+\hat{\mathbf{f}}_{j}(\mathbf{x})}{2}-\hat{\mathbf{K}}_{i,j}(\mathbf{x})\geq\Delta_{L}\sqrt{\frac{\log(1/\delta)}{8}\left(n_{i}^{-1}+n_{j}^{-1}\right)}\right)\leq\delta.$ (10)
Note that the event in (9) does not depend on $n_{i^{\prime}}$ for $i^{\prime}\neq i$, and the event in (10) does not depend on $n_{i^{\prime}}$ for $i^{\prime}\notin\{i,j\}$. By union bound, all the events in (9) and (10) do not hold for all $i\leq j$ and $n_{1},\ldots,n_{m}\in[T]$ with probability at least
$1-mT\delta-\frac{m(m-1)}{2}T^{2}\delta\geq 1-\frac{m^{2}}{2}T^{2}\delta.$
If these events do not hold, then
$L(\boldsymbol{\alpha})-\hat{L}(\boldsymbol{\alpha};\mathbf{x})$
$=\sum_{i,j}\alpha_{i}\alpha_{j}\left(\frac{\mathbf{f}_{i}+\mathbf{f}_{j}}{2}+\mathbf{K}_{i,j}-\frac{\hat{\mathbf{f}}_{i}(\mathbf{x})+\hat{\mathbf{f}}_{j}(\mathbf{x})}{2}-\hat{\mathbf{K}}_{i,j}(\mathbf{x})\right)$
$\leq\sum_{i}\alpha_{i}^{2}\left(\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\right)+\sum_{(i,j)\in[m]^{2},i\neq j}\alpha_{i}\alpha_{j}\Delta_{L}\sqrt{\frac{\log(1/\delta)}{8}\left(n_{i}^{-1}+n_{j}^{-1}\right)}$
$\leq\sum_{i}\alpha_{i}^{2}\left(\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\right)+\Delta_{L}\sqrt{\frac{\log(1/\delta)}{8}}\sum_{(i,j)\in[m]^{2},i\neq j}\alpha_{i}\alpha_{j}\left(n_{i}^{-1/2}+n_{j}^{-1/2}\right)$
$=\sum_{i}\alpha_{i}^{2}\frac{\Delta_{\kappa}}{n_{i}}+\Delta_{L}\sqrt{\frac{\log(1/\delta)}{8}}\sum_{(i,j)\in[m]^{2}}\alpha_{i}\alpha_{j}\left(n_{i}^{-1/2}+n_{j}^{-1/2}\right)$
$=\sum_{i}\alpha_{i}^{2}\frac{\Delta_{\kappa}}{n_{i}}+\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2}}\sum_{i}\alpha_{i}n_{i}^{-1/2}$
$\leq\sum_{i}\left(\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\right)\alpha_{i}.$
The other direction of the inequality is similar.
We finally prove Theorem 2.
Proof: Assume $m\geq 2$, $\beta\geq 4$ and $T\geq 2$. If $T\leq 40m$, then since $T^{-1}\log T$ is decreasing for $T\geq 3$ (the following inequalities are obviously true for $T=2$),
$\sqrt{\frac{\beta m\log T}{T}}\geq\sqrt{\frac{4m\log(40m)}{40m}}\geq\sqrt{\frac{\log 80}{10}}\geq 0.66,$
and Theorem 2 is trivially true since $\mathbb{E}\big{[}L(\hat{P}^{(T)})\big{]}-\min_{\boldsymbol{\alpha}}L(\boldsymbol{\alpha})\leq\Delta_{L}$. Hence we can assume $T\geq 40m+1$.
Let $\boldsymbol{\alpha}^{*}$ be the minimizer of $L(\boldsymbol{\alpha})$. Let $\bar{x}^{(t)}$ be the sample obtained at the $t$-th pull. Let $\alpha_{i}^{(t-1)}=\mathbf{1}\{t=i\}$ for $t\in[m]$, so “$\bar{x}^{(t)}$ is generated from the distribution $P_{i}$ with probability $\alpha_{i}^{(t-1)}$” holds for every $t\geq 1$. Write $\bar{x}^{([s])}:=(\bar{x}^{(t)})_{t\in[s]}$. For $s<t$, let $\hat{x}^{(s)}$ be a random variable with the same conditional distribution given $\bar{x}^{([s-1])}$ as $\bar{x}^{(s)}$, but is conditionally independent of all other random variables given $\bar{x}^{([s-1])}$. The joint distribution of $\bar{x}^{([t-1])},\bar{x}^{(t)},\hat{x}^{(s)}$ is
$P_{\bar{x}^{([t-1])},\bar{x}^{(t)},\hat{x}^{(s)}}=P_{\bar{x}^{([t-1])},\bar{x}^{(t)}}P_{\bar{x}^{(s)}|\bar{x}^{([s-1])}}.$
Recall that $\bar{x}^{(s)}$ is generated from the distribution $P_{i}$ with probability $\alpha_{i}^{(s-1)}$ for $i\in[m]$, where $\boldsymbol{\alpha}^{(s-1)}=(\alpha_{i}^{(s-1)})_{i\in[m]}$ is computed using $\bar{x}^{([s-1])}$. We have
$\mathbb{E}[\kappa(\hat{x}^{(s)},\bar{x}^{(t)})]$
$=\mathbb{E}\left[\mathbb{E}\left[\kappa(\hat{x}^{(s)},\bar{x}^{(t)})\,\big{|}\,\bar{x}^{([t-1])}\right]\right]$
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}\left[\sum_{i=1}^{m}\sum_{j=1}^{m}\alpha_{i}^{(s-1)}\alpha_{i}^{(t-1)}\mathbb{E}_{X\sim P_{i},X^{\prime}\sim P_{j}}\left[\kappa(X,X^{\prime})\right]\right]$
$=\mathbb{E}\left[(\boldsymbol{\alpha}^{(s-1)})^{\top}\mathbf{K}\boldsymbol{\alpha}^{(t-1)}\right],$
where (a) is because $\hat{x}^{(s)}$ only depends on $\bar{x}^{([s-1])}$ (i.e., is conditionally independent of all other random variables in the expression given $\bar{x}^{([s-1])}$), and $\bar{x}^{(t)}$ only depends on $\bar{x}^{([t-1])}$. Write $\delta_{\mathrm{TV}}(A\|B)$ for the total variation distance between the distributions of the random variables $A$ and $B$. Write $I(A;B|C)$ for the conditional mutual information between $A$ and $B$ given $C$ in nats. We have
$\mathbb{E}[\kappa(\bar{x}^{(s)},\bar{x}^{(t)})]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\mathbb{E}[\kappa(\hat{x}^{(s)},\bar{x}^{(t)})]+\Delta_{\kappa}\delta_{\mathrm{TV}}(\bar{x}^{(s)},\bar{x}^{(t)}\,\|\,\hat{x}^{(s)},\bar{x}^{(t)})$
$\leq\mathbb{E}\left[(\boldsymbol{\alpha}^{(s-1)})^{\top}\mathbf{K}\boldsymbol{\alpha}^{(t-1)}\right]+\Delta_{\kappa}\delta_{\mathrm{TV}}(\bar{x}^{([s-1])},\bar{x}^{(s)},\bar{x}^{(t)}\,\|\,\bar{x}^{([s-1])},\hat{x}^{(s)},\bar{x}^{(t)})$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\mathbb{E}\left[(\boldsymbol{\alpha}^{(s-1)})^{\top}\mathbf{K}\boldsymbol{\alpha}^{(t-1)}\right]+\Delta_{\kappa}\sqrt{\frac{1}{2}I(\bar{x}^{(s)};\bar{x}^{(t)}|\bar{x}^{([s-1])})},$
where (b) is because $\kappa$ takes values over $[\kappa_{0},\kappa_{1}]$ with $\Delta_{\kappa}=\kappa_{1}-\kappa_{0}$, and (c) is by Pinsker’s inequality. We also have, for every $t$,
$\mathbb{E}[\kappa(\bar{x}^{(t)},\bar{x}^{(t)})]$ $\leq\mathbb{E}\left[(\boldsymbol{\alpha}^{(t-1)})^{\top}\mathbf{K}\boldsymbol{\alpha}^{(t-1)}\right]+\Delta_{\kappa}.$
Hence,
$\mathbb{E}\left[\frac{1}{T^{2}}\sum_{s=1}^{T}\sum_{t=1}^{T}\kappa(\bar{x}^{(s)},\bar{x}^{(t)})\right]-\mathbb{E}\left[\frac{1}{T^{2}}\sum_{s=1}^{T}\sum_{t=1}^{T}(\boldsymbol{\alpha}^{(s-1)})^{\top}\mathbf{K}\boldsymbol{\alpha}^{(t-1)}\right]$
$\leq\frac{2\Delta_{\kappa}}{T^{2}}\sum_{s=1}^{T}\sum_{t=s+1}^{T}\sqrt{\frac{1}{2}I(\bar{x}^{(s)};\bar{x}^{(t)}|\bar{x}^{([s-1])})}+\frac{\Delta_{\kappa}}{T}$
$=\frac{2\Delta_{\kappa}}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{t-1}\sqrt{\frac{1}{2}I(\bar{x}^{(s)};\bar{x}^{(t)}|\bar{x}^{([s-1])})}+\frac{\Delta_{\kappa}}{T}$
$\leq\frac{2\Delta_{\kappa}}{T^{2}}\sum_{t=1}^{T}\sqrt{\frac{t-1}{2}\sum_{s=1}^{t-1}I(\bar{x}^{(s)};\bar{x}^{(t)}|\bar{x}^{([s-1])})}+\frac{\Delta_{\kappa}}{T}$
=$\stackrel{{\scriptstyle(d)}}{{=}}\frac{2\Delta_{\kappa}}{T^{2}}\sum_{t=1}^{T}\sqrt{\frac{t-1}{2}I(\bar{x}^{([t-1])};\bar{x}^{(t)})}+\frac{\Delta_{\kappa}}{T}$
$\stackrel{{\scriptstyle(e)}}{{\leq}}\frac{2\Delta_{\kappa}}{T^{2}}\sum_{t=1}^{T}\sqrt{\frac{t-1}{2}\log m}+\frac{\Delta_{\kappa}}{T}$
$\leq\frac{2\Delta_{\kappa}}{T^{2}}\sqrt{\frac{\log m}{2}}\int_{0}^{T}\sqrt{\tau}\mathrm{d}\tau+\frac{\Delta_{\kappa}}{T}$
$=\frac{4\Delta_{\kappa}}{3}\sqrt{\frac{\log m}{2T}}+\frac{\Delta_{\kappa}}{T},$
where (d) is by the chain rule of mutual information, and (e) is because $\bar{x}^{(t)}$ only depends on $\bar{x}^{([t-1])}$ through the choice of arm $b^{(t)}\in[m]$, and hence $I(\bar{x}^{([t-1])};\bar{x}^{(t)})$ is upper bounded by the entropy of $b^{(t)}$, which is at most $\log m$. Also note that $\mathbb{E}[f(\bar{x}^{(t)})]=\mathbf{f}^{\top}\mathbb{E}[\boldsymbol{\alpha}^{(t-1)}]$. Hence,
$\mathbb{E}[L(\hat{P}^{(T)})]$
$=\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}f(\bar{x}^{(t)})+\frac{1}{T^{2}}\sum_{s=1}^{T}\sum_{t=1}^{T}\kappa(\bar{x}^{(s)},\bar{x}^{(t)})\right]$
$\leq\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbf{f}^{\top}\boldsymbol{\alpha}^{(t-1)}+\frac{1}{T^{2}}\sum_{s=1}^{T}\sum_{t=1}^{T}(\boldsymbol{\alpha}^{(s-1)})^{\top}\mathbf{K}\boldsymbol{\alpha}^{(t-1)}\right]+\frac{4\Delta_{\kappa}}{3}\sqrt{\frac{\log m}{2T}}+\frac{\Delta_{\kappa}}{T}$
$=\mathbb{E}\left[\mathbf{f}^{\top}\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{\alpha}^{(t-1)}+\left(\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{\alpha}^{(t-1)}\right)^{\top}\mathbf{K}\left(\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{\alpha}^{(t-1)}\right)\right]+\frac{4\Delta_{\kappa}}{3}\sqrt{\frac{\log m}{2T}}+\frac{\Delta_{\kappa}}{T}$
$=\mathbb{E}\left[L\left(\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{\alpha}^{(t-1)}\right)\right]+\frac{4\Delta_{\kappa}}{3}\sqrt{\frac{\log m}{2T}}+\frac{\Delta_{\kappa}}{T}$
$\stackrel{{\scriptstyle(f)}}{{\leq}}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t-1)})\right]+\frac{4\Delta_{\kappa}}{3}\sqrt{\frac{\log m}{2T}}+\frac{\Delta_{\kappa}}{T},$ (11)
where (f) is because $\mathbf{K}$ is positive semidefinite, and hence $L$ is convex. Therefore, to bound the optimality gap, we study the expected loss $\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t)})\right]$ of the estimate of the optimal mixture distribution $\boldsymbol{\alpha}^{(t)}$.
Let $\tilde{\delta}>0$. Let $\tilde{E}$ be the event
$L(\boldsymbol{\alpha})-\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(n_{i})_{i}})\leq\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{1}{2n_{i}}\log\frac{m^{2}T^{2}}{2\tilde{\delta}}}+\frac{\Delta_{\kappa}}{n_{i}}\right)\alpha_{i}$
for every $n_{1},\ldots,n_{m}\in[T]$ and probability vector $\boldsymbol{\alpha}$, as in Lemma 1. By Lemma 1, $\mathbb{P}(\tilde{E})\geq 1-\tilde{\delta}$.
Fix a time $t\in\{m,\ldots,T\}$. Let $E_{t}$ be the event
$\hat{L}(\boldsymbol{\alpha}^{*};\mathbf{x}^{(n_{i})_{i}})-L(\boldsymbol{\alpha}^{*})\leq\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{\beta\log t}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\right)\alpha_{i}^{*}$
for every $n_{1},\ldots,n_{m}\geq 1$ such that $\sum_{i}n_{i}=t$. Since
$\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{1}{2n_{i}}\log\frac{m^{2}t^{2}}{2m^{2}t^{-2}/2}}+\frac{\Delta_{\kappa}}{n_{i}}\right)\alpha_{i}^{*}$
$=\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{1}{2n_{i}}\log t^{4}}+\frac{\Delta_{\kappa}}{n_{i}}\right)\alpha_{i}^{*}$
$\leq\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{\beta\log t}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\right)\alpha_{i}^{*}$
by $\beta\geq 4$, applying Lemma 1,
$\mathbb{P}(E_{t})\geq 1-m^{2}t^{-2}/2.$ (12)
If the event $E_{t}$ holds, by taking $n_{i}=n_{i}^{(t)}$,
$\hat{L}(\boldsymbol{\alpha}^{*};\mathbf{x}^{(t)})-L(\boldsymbol{\alpha}^{*})\leq(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}^{*}.$ (13)
If the event $\tilde{E}$ holds,
$L(\boldsymbol{\alpha})-\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})$
$\leq\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{1}{2n_{i}^{(t)}}\log\frac{m^{2}T^{2}}{2\tilde{\delta}}}+\frac{\Delta_{\kappa}}{n_{i}^{(t)}}\right)\alpha_{i}$ (14)
for every $\boldsymbol{\alpha}$. Combining (13) and (14) (with $\boldsymbol{\alpha}=\boldsymbol{\alpha}^{(t)}$),
$\hat{L}(\boldsymbol{\alpha}^{(t)};\mathbf{x}^{(t)})-\hat{L}(\boldsymbol{\alpha}^{*};\mathbf{x}^{(t)})+(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}^{*}$
$+\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{1}{2n_{i}^{(t)}}\log\frac{m^{2}T^{2}}{2\tilde{\delta}}}+\frac{\Delta_{\kappa}}{n_{i}^{(t)}}\right)\alpha_{i}^{(t)}$
$\geq L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*}).$
By (6), $\hat{L}(\boldsymbol{\alpha}^{(t)};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}^{(t)}\leq\hat{L}(\boldsymbol{\alpha}^{*};\mathbf{x}^{(t)})-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}^{*}$, and hence if the events $\tilde{E},E_{t}$ hold,
$(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}^{(t)}+\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{1}{2n_{i}^{(t)}}\log\frac{m^{2}T^{2}}{2\tilde{\delta}}}+\frac{\Delta_{\kappa}}{n_{i}^{(t)}}\right)\alpha_{i}^{(t)}$
$\geq L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*}).$ (15)
We have
$(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}^{(t)}+\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{1}{2n_{i}^{(t)}}\log\frac{m^{2}T^{2}}{2\tilde{\delta}}}+\frac{\Delta_{\kappa}}{n_{i}^{(t)}}\right)\alpha_{i}^{(t)}$
$=\sum_{i=1}^{m}\left(\Delta_{L}\sqrt{\frac{\beta\log t}{2n_{i}^{(t)}}}+\frac{\Delta_{\kappa}}{n_{i}^{(t)}}+\Delta_{L}\sqrt{\frac{1}{2n_{i}^{(t)}}\log\frac{m^{2}T^{2}}{2\tilde{\delta}}}+\frac{\Delta_{\kappa}}{n_{i}^{(t)}}\right)\alpha_{i}^{(t)}$
$=\sum_{i}\left(\frac{\Delta_{L}}{\sqrt{n_{i}^{(t)}}}\left(\sqrt{\frac{\beta}{2}\log t}+\sqrt{\frac{1}{2}\log\frac{m^{2}T^{2}}{2\tilde{\delta}}}\right)+\frac{2\Delta_{\kappa}}{n_{i}^{(t)}}\right)\alpha_{i}^{(t)}$
$\leq\sum_{i}\left(\frac{\Delta_{L}}{\sqrt{n_{i}^{(t)}}}\left(\sqrt{\frac{\beta}{2}\log T}+\sqrt{\frac{1}{2}\log\frac{m^{2}T^{2}}{2\tilde{\delta}}}\right)+\frac{\Delta_{L}}{\sqrt{n_{i}^{(t)}}}\right)\alpha_{i}^{(t)}$
$=\Delta_{L}\eta\sum_{i}\frac{\alpha_{i}^{(t)}}{\sqrt{n_{i}^{(t)}}},$
where
$\eta:=\sqrt{\frac{\beta}{2}\log T}+\sqrt{\frac{1}{2}\log\frac{m^{2}T^{2}}{2\tilde{\delta}}}+1.$
Substituting into (15), if the events $\tilde{E},E_{t}$ hold,
$\Delta_{L}\eta\sum_{i}\frac{\alpha_{i}^{(t)}}{\sqrt{n_{i}^{(t)}}}\geq L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*}).$
Hence, in general (regardless of whether $\tilde{E},E_{t}$ hold), denoting the indicator function of $\tilde{E}\cap E_{t}$ as $\mathbf{1}_{\tilde{E}\cap E_{t}}\in\{0,1\}$,
$\sum_{i}\frac{\alpha_{i}^{(t)}}{\sqrt{n_{i}^{(t)}}}\geq\frac{L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})}{\Delta_{L}\eta}\mathbf{1}_{\tilde{E}\cap E_{t}}.$
Let
$\Psi^{(t)}:=\sum_{i=1}^{m}\psi(n_{i}^{(t)}-1),$
where $\psi(n):=\sum_{i=1}^{n}i^{-1/2}$. Recall that we pull arm $i$ at time $t+1$ with probability $\alpha_{i}^{(t)}$. The expected increase of $\Psi^{(t)}$ is
$\mathbb{E}\left[\Psi^{(t+1)}-\Psi^{(t)}\,\big{|}\,\mathbf{x}^{(t)}\right]$ $=\sum_{i=1}^{m}\left(\psi(n_{i}^{(t)})-\psi(n_{i}^{(t)}-1)\right)\alpha_{i}^{(t)}$
$=\sum_{i=1}^{m}\frac{\alpha_{i}^{(t)}}{\sqrt{n_{i}^{(t)}}}$
$\geq\frac{L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})}{\Delta_{L}\eta}\mathbf{1}_{\tilde{E}\cap E_{t}}.$
Note that
$\Psi^{(T)}$ $=\sum_{i=1}^{m}\psi(n_{i}^{(T)}-1)$
$\leq\sum_{i=1}^{m}\int_{0}^{n_{i}^{(T)}-1}\min\{\tau^{-1/2},1\}\mathrm{d}\tau$
$\stackrel{{\scriptstyle(a)}}{{\leq}}m\int_{0}^{m^{-1}\sum_{i=1}^{m}n_{i}^{(T)}-1}\min\{\tau^{-1/2},1\}\mathrm{d}\tau$
$=m\int_{0}^{T/m-1}\min\{\tau^{-1/2},1\}\mathrm{d}\tau$
=$\stackrel{{\scriptstyle(b)}}{{=}}m\left(1+\int_{1}^{T/m-1}\tau^{-1/2}\mathrm{d}\tau\right)$
$=m\left(2\sqrt{\frac{T}{m}-1}-1\right),$
where (a) is because $a\mapsto\int_{0}^{a}\min\{\tau^{-1/2},1\}\mathrm{d}\tau$ is concave, and (b) is because $T\geq 40m+1$, so $T/m-1\geq 1$. Therefore,
$m\left(2\sqrt{\frac{T}{m}-1}-1\right)$ $\geq\mathbb{E}\left[\Psi^{(T)}-\Psi^{(m)}\right]$
$=\sum_{t=m}^{T-1}\mathbb{E}\left[\Psi^{(t+1)}-\Psi^{(t)}\right]$
$\geq\sum_{t=m}^{T-1}\mathbb{E}\left[\frac{L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})}{\Delta_{L}\eta}\mathbf{1}_{\tilde{E}\cap E_{t}}\right]$
$\geq\frac{1}{\Delta_{L}\eta}\sum_{t=m}^{T-1}\left(\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})\right]-\Delta_{L}\mathbb{P}((\tilde{E}\cap E_{t})^{c})\right)$
$\stackrel{{\scriptstyle(c)}}{{\geq}}\frac{1}{\Delta_{L}\eta}\sum_{t=m}^{T-1}\left(\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})\right]-\Delta_{L}(\tilde{\delta}+m^{2}t^{-2}/2)\right)$
$\geq\frac{1}{\Delta_{L}\eta}\sum_{t=m}^{T-1}\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})\right]-\frac{T\tilde{\delta}}{\eta}-\frac{m^{2}}{2\eta}\int_{m-1}^{T-1}t^{-2}\mathrm{d}t$
$\geq\frac{1}{\Delta_{L}\eta}\sum_{t=m}^{T-1}\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})\right]-\frac{T\tilde{\delta}}{\eta}-\frac{m^{2}}{2\eta(m-1)}$
$\geq\frac{1}{\Delta_{L}\eta}\sum_{t=m}^{T-1}\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})\right]-\frac{T\tilde{\delta}}{\eta}-\frac{m}{\eta},$
where (c) is by (12). Hence,
$\frac{1}{\Delta_{L}}\sum_{t=0}^{T-1}\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})\right]$
$\leq\frac{1}{\Delta_{L}}\sum_{t=m}^{T-1}\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t)})-L(\boldsymbol{\alpha}^{*})\right]+m$
$\leq\eta m\left(2\sqrt{\frac{T}{m}-1}-1\right)+T\tilde{\delta}+2m$
=$\stackrel{{\scriptstyle(d)}}{{=}}m\left(2\sqrt{\frac{T}{m}-1}-1\right)\left(\sqrt{\frac{\beta}{2}\log T}+\sqrt{\frac{1}{2}\log\frac{mT^{3}}{2}}+1\right)+3m$
$\stackrel{{\scriptstyle(e)}}{{\leq}}m\left(2\sqrt{\frac{T}{m}-1}\right)\left(\sqrt{\frac{\beta}{2}\log T}+\sqrt{\frac{1}{2}\log\frac{mT^{3}}{2}}+1\right)$
$\leq 2\sqrt{mT}\left(\sqrt{\frac{\beta}{2}\log T}+\sqrt{\frac{1}{2}\log\frac{mT^{3}}{2}}+1\right),$
where (d) is by substituting $\tilde{\delta}=m/T$, (e) is because $\sqrt{\frac{\beta}{2}\log T}\geq\sqrt{2\log 81}\geq 2.9$ and $\sqrt{\frac{1}{2}\log\frac{mT^{3}}{2}}\geq 2.5$ (recall that $T\geq 40m+1\geq 81$). Substituting into (11),
$\frac{1}{\Delta_{L}}\left(\mathbb{E}[L(\hat{P}^{(T)})]-L(\boldsymbol{\alpha}^{*})\right)$
$\leq\frac{1}{\Delta_{L}T}\sum_{t=1}^{T}\mathbb{E}\left[L(\boldsymbol{\alpha}^{(t-1)})-L(\boldsymbol{\alpha}^{*})\right]+\frac{4\Delta_{\kappa}}{3\Delta_{L}}\sqrt{\frac{\log m}{2T}}+\frac{\Delta_{\kappa}}{T\Delta_{L}}$
$\leq 2\sqrt{\frac{m}{T}}\left(\sqrt{\frac{\beta}{2}\log T}+\sqrt{\frac{1}{2}\log\frac{mT^{3}}{2}}+1\right)+\frac{2}{3}\sqrt{\frac{\log m}{2T}}+\frac{1}{2T}$
$=\frac{1}{\sqrt{T}}\left(2\sqrt{\frac{\beta}{2}m\log T}+\sqrt{2m\log\frac{mT^{3}}{2}}+2\sqrt{m}+\frac{2}{3}\sqrt{\frac{\log m}{2}}+\frac{1}{2\sqrt{T}}\right)$
$\leq\frac{1}{\sqrt{T}}\bigg{(}2\sqrt{\frac{\beta}{2}m\log T}+\sqrt{2m\log T^{4}}+\frac{2\sqrt{m\log T}}{\sqrt{\log 81}}$
$\quad\quad+\frac{2}{3}\frac{\sqrt{m\log T}}{\sqrt{m\log(40m+1)}}\sqrt{\frac{\log m}{2}}+\frac{1}{2\sqrt{81}}\cdot\frac{\sqrt{m\log T}}{\sqrt{2\log 81}}\bigg{)}$
$=\sqrt{\frac{m\log T}{T}}\left(\sqrt{2\beta}+2\sqrt{2}+\frac{2}{\sqrt{\log 81}}+\frac{2}{3}\sqrt{\frac{\log m}{2m\log(40m+1)}}+\frac{1}{2\sqrt{81}}\cdot\frac{1}{\sqrt{2\log 81}}\right)$
$\leq\sqrt{\frac{m\log T}{T}}\left(\sqrt{2\beta}+2\sqrt{2}+\frac{2}{\sqrt{\log 81}}+\frac{2}{3}\sqrt{\frac{\log 2}{4\log 81}}+\frac{1}{2\sqrt{81}}\cdot\frac{1}{\sqrt{2\log 81}}\right)$
$\leq\sqrt{\frac{m\log T}{T}}\left(\sqrt{2\beta}+3.934\right)$
$\leq\sqrt{\frac{m\log T}{T}}\left(\sqrt{2\beta}+\frac{3.934}{2}\sqrt{\beta}\right)$
$\leq 3.382\sqrt{\frac{\beta m\log T}{T}}.$
This completes the proof of Theorem 2 (with an improved constant $3.382$ instead of $4$).
Appendix B Sparse Mixture Upper Confidence Bound-Continuum-Armed Bandit Algorithm
Algorithm 3 Sparse-Mixture-UCB-CAB
1:Input: $m$ generative arms, number of rounds $T$
2:Output: Gathered samples $\mathbf{x}^{(T)}$
3:Initialize the set of subscribed arms $\mathcal{S}\leftarrow[m]$.
4:for $t\in\{0,\ldots,m-1\}$  do
5:     Pull arm $t+1$ at time $t+1$ to obtain sample $x_{t+1,1}\sim P_{t+1}$. Set $n_{t+1}^{(m)}=1$.
6:end for
7:for $t\in\{m,\ldots,T-1\}$  do
8:     repeat
9:         Compute
$\boldsymbol{\alpha}^{(t)}\leftarrow\underset{\boldsymbol{\alpha}:\,\mathrm{supp}(\boldsymbol{\alpha})\subseteq\mathcal{S}}{\mathrm{argmin}}\left(\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})+\lambda|\mathcal{S}|-(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}\right),$ (16)
where $\boldsymbol{\epsilon}^{(t)}\in\mathbb{R}^{m}$ is defined in (7). Let the minimum value above be $C$.
10:         Compute the following “worst arm” if $|\mathcal{S}|\geq 2$:
$i^{\prime}\leftarrow\underset{i\in\mathcal{S}}{\mathrm{argmin}}\min_{\boldsymbol{\alpha}:\,\mathrm{supp}(\boldsymbol{\alpha})\subseteq\mathcal{S}\backslash\{i\}}\left(\hat{L}(\boldsymbol{\alpha};\mathbf{x}^{(t)})+\lambda(|\mathcal{S}|-1)+(\boldsymbol{\epsilon}^{(t)})^{\top}\boldsymbol{\alpha}\right).$
Let the minimum value above be $C^{\prime}$.
11:         if $C^{\prime}\leq C$ then
12:              Unsubscribe arm $i^{\prime}$ (i.e., $\mathcal{S}\leftarrow\mathcal{S}\backslash\{i^{\prime}\}$)
13:         end if
14:     until no more arms are unsubscribed
15:     Generate the arm index $b^{(t+1)}\in[m]$ at random with $\mathbb{P}(b^{(t+1)}=i)=\alpha_{i}^{(t)}$.
16:     Pull arm $b=b^{(t+1)}$ at time $t+1$ to obtain a new sample $x_{b,n_{b}^{(t)}+1}\sim P_{b}$. Set $n_{b}^{(t+1)}=n_{b}^{(t)}+1$ and $n_{j}^{(t+1)}=n_{j}^{(t)}$ for $j\neq b$.
17:end forreturn samples $\mathbf{x}^{(T)}$
Appendix C Additional Numerical Results
C.1 Details of the Numerical Experiments
Hyper-parameter Choice.
The kernel bandwidths for the RKE and KID metrics were chosen based on the guidelines provided in their respective papers to ensure clear distinction between models. The values for $\Delta_{L}$ and $\Delta_{\kappa}$ in our online algorithms (7) were set according to the magnitudes of the metrics and their behavior on a validation subset. The number of sampling rounds was adjusted according to the number of models and metric convergence, both of which depend on the bandwidth. To ensure the statistical significance of results, all experiments were repeated 10 times with different random seeds, and the reported plots represent the average results.
C.2 Optimal Mixture for Quality and Diversity via KID
Suppose $P$ is the distribution of generated images of a model, and $Q$ is the target distribution. Recall that for KID (3), we take the quadratic term to be $\kappa(x,x^{\prime})=k(\psi(x),\psi(x^{\prime}))$ (with an expectation $\mathbb{E}_{X,X^{\prime}\sim P}[k(\psi(X),\psi(X^{\prime}))])$) and the linear term to be $f(x)=-2\mathbb{E}_{Y\sim Q}[k(\psi(x),\psi(Y))]$ (with an expectation $-2\mathbb{E}_{X\sim P,Y\sim Q}[k(\psi(X),\psi(Y))]$). In order to run our online algorithms, we use $\Delta_{L}$ and $\Delta_{\kappa}$ based on a validation portion to make sure the UCB terms have the right magnitude for forcing exploration.
FFHQ Generated Images.
In this experiment, we use images generated by five different models: LDM [36], StyleGAN-XL [37], Efficient-vdVAE [38], Insgen [39], and StyleNAT [40]. We use 10,000 images from each model and a kernel bandwidth of 40 to determine the optimal mixture, resulting in the weights (0.33, 0.57, 0, 0, 0.10). The online algorithms run for 8,000 sampling rounds. The quality and diversity scores for each model, including the results for the optimal mixture based on KID, are presented in Table 1.
In Tables 1 and 2, we observe that the Precision of the optimal mixture is similar to that of the maximum Precision score among individual models. On the other hand, the Recall-based diversity improved in the mixture case. However, the quality-measuring Density score slightly decreased for the selected mixture model, as Density is a linear score for quality that could be optimized by an individual model. On the other hand, the Coverage score of the mixture model was higher than each individual model.
Note that Precision and Density are scores on the average quality of samples. Intuitively, the quality score of a mixture of models is the average of the quality score of the individual models, and hence the quality score of a mixture cannot be better than the best individual model. On the other hand, Recall and Coverage measure the diversity of the samples, which can increase by considering a mixture of the models. To evaluate the net diversity-quality effect, we measured the FID score of the selected mixture and the best individual model, and the selected mixture model had a better FID score compared to the individual model with the best FID.
Model Precision $\uparrow$ Recall $\uparrow$ Density $\uparrow$ Coverage $\uparrow$ FID $\downarrow$
LDM 0.856 ± 0.008 0.482 ± 0.008 0.959 ± 0.027 0.776 ± 0.006 189.876 ± 1.976
StyleGAN-XL 0.798 ± 0.007 0.515 ± 0.007 0.726 ± 0.018 0.691 ± 0.009 186.163 ± 2.752
Efficient-vdVAE 0.854 ± 0.011 0.143 ± 0.007 0.952 ± 0.033 0.545 ± 0.008 490.385 ± 4.377
Insgen 0.76 ± 0.006 0.281 ± 0.007 0.716 ± 0.016 0.692 ± 0.005 278.235 ± 1.617
StyleNAT 0.834 ± 0.008 0.478 ± 0.007 0.867 ± 0.023 0.775 ± 0.007 185.067 ± 2.123
Optimal Mixture (KID) 0.818 ± 0.007 0.57 ± 0.008 0.816 ± 0.025 0.765 ± 0.007 168.127 ± 1.596
Mixture-UCB-CAB (KID) 0.828 ± 0.007 0.571 ± 0.008 0.838 ± 0.016 0.787 ± 0.007 170.578 ± 2.075
Mixture-UCB-OGD (KID) 0.827 ± 0.006 0.573 ± 0.005 0.825 ± 0.015 0.787 ± 0.006 170.113 ± 1.930
Table 1: Quality and diversity scores for the FFHQ experiment, including precision, recall, density, coverage, and FID metrics ($\pm$ standard deviation).
LSUN-Bedroom
We use images generated by four different models: StyleGAN [41], Projected GAN [42], iDDPM [43], and Unleashing Transformers [44]. We utilize 10,000 images from each model to calculate the optimal mixture, resulting in weights of (0.51, 0, 0.49, 0). A kernel bandwidth of 40 is used and the algorithms run for 8,000 sampling steps. The quality and diversity scores for each model, including the results for the optimal mixture based on KID, are presented in Table 2.
Model Precision $\uparrow$ Recall $\uparrow$ Density $\uparrow$ Coverage $\uparrow$ FID $\downarrow$
StyleGAN 0.838 ± 0.008 0.446 ± 0.007 0.941 ± 0.019 0.821 ± 0.004 175.575 ± 2.055
Projected GAN 0.749 ± 0.015 0.329 ± 0.008 0.592 ± 0.027 0.517 ± 0.008 324.066 ± 3.753
iDDPM 0.838 ± 0.006 0.641 ± 0.006 0.660 ± 0.018 0.825 ± 0.006 154.680 ± 3.036
Unleashing Transformers 0.786 ± 0.008 0.449 ± 0.006 0.649 ± 0.013 0.581 ± 0.013 339.982 ± 6.118
Optimal Mixture (KID) 0.838 ± 0.006 0.589 ± 0.005 0.900 ± 0.016 0.833 ± 0.004 149.779 ± 2.238
Mixture-UCB-CAB (KID) 0.835 ± 0.007 0.602 ± 0.007 0.894 ± 0.014 0.834 ± 0.005 151.28 ± 1.801
Mixture-UCB-OGD (KID) 0.838 ± 0.008 0.599 ± 0.007 0.902 ± 0.023 0.834 ± 0.006 151.10 ± 1.906
Table 2: Quality and diversity scores for the LSUN-Bedroom experiment, including precision, recall, density, coverage, and FID metrics ($\pm$ standard deviation).
Truncated FFHQ.
We use StyleGAN2-ADA [47] trained on FFHQ dataset to generate images. We randomly choose 8 initial points and use the Truncation Method [46, 41] to generate images with limited diversity around each of the chosen points. We use truncation value of 0.3 and generate 5000 images from each model to find the optimal mixture. The optimal weights for the mixture is (0.07, 0.28, 0.10, 0.04, 0.21, 0.11, 0.12, 0.07). A kernel bandwidth of 40 is used and 4,000 sampling steps are conducted.
C.3 Optimal Mixture for Diversity via RKE
Truncated FFHQ.
We employ StyleGAN2-ADA [47], trained on the FFHQ dataset, to generate images. Eight initial points are randomly selected, and the Truncation Method [46, 41] is applied with a truncation value of 0.3 to generate images with limited diversity around these points. For the offline optimization, 5,000 images are generated from each model, using a kernel bandwidth of 40 to identify the optimal mixture. In the online experiment, sampling is conducted over 2,000 steps.
Figure 8: Visual demonstration of the increase in diversity when mixing arms compared to individual arms for truncated FFHQ generative models. The RKE values for each model and the mixture serve as indicators of diversity.
Truncated AFHQ Cat.
Similar to the previous experiment, we use StyleGAN2-ADA to generate AFHQ Cat images. Four initial points are selected, and a truncation value of 0.6 is applied to simulate diversity-controlled models. For the quadratic optimization, 5,000 images are generated from each model, with sampling over 1,200 steps to determine the optimal mixture.
Figure 9: Visual demonstration of the increase in diversity when mixing arms compared to individual arms for truncated AFHQ Cat generative models. The RKE values for each model and the mixture represent the diversity.
Style-Specific Generators.
We use Stable Diffusion XL to generate images of cars in distinct styles: realistic, surreal, and cartoon. For this experiment, we utilize 2,000 images from each model to determine the optimal mixture, which yields weights of (0.67, 0.27, 0.06). This mixture increases the RKE value from 7.86 (the optimal value of the realistic images) to 9.19. We set the kernel bandwidth to 30 and execute the online algorithms over 1,000 sampling steps.
Sofa Images.
We generate images of the object “Sofa” using prompts with descriptions of the environment across the models FLUX.1-Schnell [50], Kandinsky 3.0 [51], PixArt-$\alpha$ [52], and Stable Diffusion XL [49]. Solving the RKE optimization with 1,000 images reveals that sampling 38% from FLUX and 62% from Kandinsky improves the RKE score from the one-arm optimum of 12.12 to 14.51. We set the kernel bandwidth to 30 and conduct the online experiment over 700 steps. We observe that Mixture-UCB-OGD achieves noticeably faster convergence to the optimal mixture RKE in this scenario.
The prompts follow the structure: “A adjective sofa is verb in a location,” with the terms for Adjective, Verb, and Location generated by GPT-4o [59], specifically for the object ”Sofa.”
Figure 10: Visual comparison of diversity of each arm and the mixture for sofa image generators
Red Bird Images.
To observe the performance of mixing the models while generating images on a single prompt, we generate images with the prompt “Red bird, cartoon style” using Kandinsky 3.0 [51], Stable Diffusion 3-medium [53], and PixArt-$\alpha$ [52]. We use 8000 images and a kernel bandwidth of 30 to find the optimal mixture in an offline manner. The increase in diversity is shown in Figure 1. We observe a noticeable boost in the RKE and Vendi scores, showing the diversity has improved. The performance of our online algorithms and the comparison of samples generated by each online algorithm are shown in Figure 2.
Dog Breeds Images.
Stable Diffusion XL is used to generate images of three breeds of dogs: Poodle, Bulldog, and German Shepherd. As shown in Figure 5, using a mixture of models results in an increase in mode count from 1.5 to 3, supporting our claim of enhanced diversity. We set the kernel bandwidth to 50 and generate 1,000 images for each breed to determine the optimal mixture, which was (0.33, 0.31, 0.36). The online algorithms are executed for 500 sampling steps.
(a) Text generators
(b) Sparse mixture
Figure 11: Comparison of online algorithms for the RKE metric on text generative models and the Sparse Mixture algorithm for FFHQ truncated generators
Text Generative Models
We utilize the OpenLLMText dataset [54], which consists of 60,000 human texts rephrased paragraph by paragraph using the models GPT2-XL [55], LLaMA-7B [56], and PaLM [57]. To extract features from each text, we employ the RoBERTa text encoder [35]. By solving the optimization problem on 10,000 texts from each model, we find that mixing the models with probabilities (0.02, 0.34, 0.64, 0) achieves the optimal mixture, improving the RKE from an optimal single model score of 69.3 to 75.2. A bandwidth of 0.6 is used for the kernel, and we run the online algorithms for 7,000 steps to demonstrate their performance.
Sparse Mixture
Four different initial points and StyleGAN2-ADA are used to generate images with a truncation of 0.6 around the points, simulating diversity-controlled models. A value of $\lambda=0.06$ and a bandwidth of 30 are selected based on the magnitudes of RKEs from the validation dataset to determine when to “unsubscribe” models in the Sparse-Mixure-UCB-CAB algorithm. We conduct three scenarios, gradually reducing the number of models to between one and three, and present a comparison of the resulting plots and their convergence values in Figure 11(b).
C.4 Optimal Mixture for Diversity and Quality via RKE and Precision/Density
Figure 12: Visual demonstration of the effect of combining Precision/Density with RKE. The CIFAR10 generator is excluded when these quality metrics are applied.
In this experiment, we use four arms: three of them are StyleGAN2-ADA models trained on FFHQ, each using a truncation value of 0.3 around a randomly selected point. The fourth arm is StyleGAN2-ADA trained on CIFAR-10. We generated 5,000 images and use a kernel bandwidth of 30 to calculate the optimal mixture. When optimizing purely for diversity using the RKE metric, the high diversity of the fourth model leads to a probability of 0.91 being assigned to it, as shown in Figure 12. However, despite the increased diversity, the quality of the generated images, based on the reference distribution, is unsatisfactory.
To address this, we incorporate a quality metric, specifically Precision/Density, into the optimization. We use the RKE score as K and the weighted Precision/Density as f according to equation 5. The weights for the quality metrics $(\lambda=0.2$ and 0.5)0.5)0.5 ) are selected based on validation data to ensure comparable scaling between the two metrics. As a result, Figure 12 shows that the fourth arm, which has low-quality outputs, is assigned a weight of zero. The online algorithms also run for 4,000 steps, with the results depicted in Figure 7.