### v2

# Be More Diverse than the Most Diverse: Optimal Mixtures of Generative Models via Mixture-UCB Bandit Algorithms

Abstract
The availability of multiple training algorithms and architectures for generative models requires a selection mechanism to form a single model over a group of well-trained generation models. The selection task is commonly addressed by identifying the model that maximizes an evaluation score based on the diversity and quality of the generated data. However, such a best-model identification approach overlooks the possibility that a mixture of available models can outperform each individual model. In this work, we numerically show that a mixture of generative models on benchmark image datasets can indeed achieve a better evaluation score (based on FID and KID scores), compared to the individual models. This observation motivates the development of efficient algorithms for selecting the optimal mixture of the models. To address this, we formulate a quadratic optimization problem to find an optimal mixture model achieving the maximum of kernel-based evaluation scores including kernel inception distance (KID) and Rényi kernel entropy (RKE). To identify the optimal mixture of the models using the fewest possible sample queries, we view the selection task as a multi-armed bandit (MAB) problem and propose the Mixture Upper Confidence Bound (Mixture-UCB) algorithm that provably converges to the optimal mixture of the involved models. More broadly, the proposed Mixture-UCB can be extended to optimize every convex quadratic function of the mixture weights in a general MAB setting. We prove a regret bound for the Mixture-UCB algorithm and perform several numerical experiments to show the success of Mixture-UCB in finding the optimal mixture of text and image generative models. The project code is available at https://github.com/Rezaei-Parham/Mixture-UCB.
1 Introduction
The rapid advancements in generative modeling have created a need for mechanisms to combine multiple well-trained generative models, each developed using different algorithms and architectures, into a single unified model. Consider $m$ unconditional generative models $\mathcal{G}_{1},\ldots,\mathcal{G}_{m}$, where each $\mathcal{G}_{i}$ represents a probability model $P_{\mathcal{G}_{i}}$ according to which new samples are generated. A common approach for creating a unified model is to compute evaluation scores (e.g., the standard FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018) scores) that quantify the diversity and fidelity of the generated data, followed by selecting the model $P_{\mathcal{G}_{i^{*}}}$ with the best evaluated score. This best-score model selection strategy has been widely adopted for choosing generative models across various domains, including image, text, and video data generation.
Figure 1: A mixture (right-most case) of FFHQ pre-trained generative models with weights (0.25,0.4,0.01,0.08,0.26) achieves better FID and KID scores compared to each of the five involved models. The mixture weights are computed using our proposed Mixture-UCB-OGD algorithm.
However, the model selection approach by identifying the score-maximizing model overlooks the possibility that a mixture of the generative models $\alpha_{1}P_{\mathcal{G}_{1}}+\cdots+\alpha_{m}P_{\mathcal{G}_{m}}$, where each sample is generated from a randomly-selected model with $\alpha_{i}$ being the probability of selecting model $\mathcal{G}_{i}$, can outperform every individual model. This motivates the following question: Can there be real-world settings where a non-degenerate mixture of some well-trained generative models obtain a better evaluation score compared to each individual model? Note that the standard FID and KID scores are convex functions of the generative model’s distribution, and thus they can be optimized by a non-degenerate mixture of the models. In this work, we numerically show that it is possible for a mixture of real-world generative models to improve evaluation scores over the individual models. An example is shown in Figure 1, where we find a non-degenerate mixture of five generative models pre-trained on the FFHQ dataset [^1] The pre-trained generative models are downloaded from dgm-eval GitHub repository (Stein et al., 2023).. As shown, assigning mixture weights $(0.25,0.4,0.01,0.08,0.26)$ to the models results in a significantly better FID score $170.11$ than the best individual FID $185.07$.
To understand the improvement achieved by the mixture model, we note that the FID and KID scores evaluate both the quality and diversity of the generated data. While the averaged quality of generated samples represents an expected value that is optimized by an individual model, the diversity of samples from a mixture of the models can significantly improve over the diversity of the individual models’ data. Figure 2 displays an illustrative example for this point, where we observe that the diversity of “red bird, cartoon style” samples generated by each of the three text-to-image models, is qualitatively and quantitatively [^2] We have evaluated the diversity scores RKE (Jalali et al., 2023) and Vendi (Dan Friedman & Dieng, 2023). lower than the diversity of their mixture. As a result, the improvement in the diversity of a mixture of generative models can result in an improved FID and KID evaluation scores, as we numerically observe in Figure 1.
1.1 Computing Optimal Mixtures of Generative Models: The Mixture-UCB Multi-Armed Bandit Algorithm
Since the evaluation score of a mixture of generative models can improve over the scores of the individual models, a natural question is how to efficiently compute the weights of an optimal mixture of the models using the fewest possible samples from the models. Here, our goal is to minimize the number of sample generation queries from sub-optimal models, which will save the time and monetary costs of identifying the best model. To achieve this, we propose viewing the task as a multi-armed bandit (MAB) problem, in which every generative model represents an arm and our goal is to find the best mixture of the models with the optimal evaluation score. The MAB approach for selecting among generative models has been recently explored by Hu et al. (2024) applying the Upper Confidence Bound (UCB) algorithm to the FID score. This MAB-based model selection extends the online model selection methods for supervised models, including the successive halving strategy (Karnin et al., 2013; Jamieson & Talwalkar, 2016; Chen & Ghosh, 2024) [^3] Jamieson & Talwalkar (2016) focused on applying successive halving on hyperparameter optimization for supervised learning, whereas Chen & Ghosh (2024) focused on generative models using maximum mean discrepancy (Gretton et al., 2012) as the score..
However, in the existing MAB algorithms, the goal is to eventually converge to a single arm with the optimal score. Successive halving (Karnin et al., 2013; Jamieson & Talwalkar, 2016; Chen & Ghosh, 2024) and the UCB algorithm developed by Hu et al. (2024) will ultimately select only one generative model after a sufficient number of iterations. However, as we discussed earlier, the evaluation scores of the generated data could be higher when the sample generation follows a mixture of models rather than a single model. This observation leads to the following task: Developing an MAB algorithm that finds the optimal mixture of the arms rather than the single best arm.
In this work, we develop a general MAB method, which we call the Mixture-UCB algorithm, that can provably find the best mixture of the arms when the evaluation score is a quadratic function of the arms’ distribution, i.e. when it represents the average of scores assigned to the pairs of drawn samples. Formulating the optimization problem for a quadratic score function results in a quadratic online convex optimization problem that can be efficiently solved using the online gradient descent algorithm. More importantly, we establish a concentration bound for the quadratic function of the mixture weights, which enables us to extend the UCB algorithm to the Mixture-UCB method for the online selection of the mixture weights.
For selecting mixtures of generative models using Mixture-UCB, we focus on evaluation scores that reduce to a quadratic function of the generative model’s distribution, including Kernel Inception Distance (KID) (Bińkowski et al., 2018), Rényi Kernel Entropy (RKE) (Jalali et al., 2023) scores, as well as the quality-measuring Precision (Sajjadi et al., 2018; Kynkäänniemi et al., 2019) and Density (Naeem et al., 2020) scores, which are linear functions of the generative distribution. Among these scores, RKE provides a reference-free entropy function for assessing the diversity of generated data, making it suitable for quantifying the variety of generated samples. Our mixture-based MAB framework can therefore be applied to find the mixture model with the maximum RKE-based diversity score. Additionally, we consider a linear combination of RKE with the Precision, Density, or KID quality scores to find a mixture of models that offers the best trade-off between quality and diversity.
Figure 2: Visual comparison of the diversity across individual arms and the optimal mixture for images generated using models Kandinsky 3, Stable Diffusion 3, and PixArt-$\alpha$ with the prompt “Red bird, cartoon style”. The mixture weights are computed via the Mixture-UCB-OGD method.
We perform several numerical experiments to test the application of our proposed Mixture-UCB approach in comparison to the Vanilla-UCB and One-Arm Oracle approaches that tend to generate samples from only one of the available generative models. Our numerical results indicate that the Mixture-UCB algorithms can generate samples with higher RKE diversity scores, and tends to generate samples from a mixture of several generative models when applied to image-based generative models. Also, we test the performance of Mixture-UCB on the KID, Precision, and Density scores, which similarly result in a higher score value for the mixture model found by the Mixture-UCB algorithm. We implement the Mixture-UCB by solving the convex optimization sub-problem at every iteration and also by applying the online gradient descent algorithm at every iteration. In our experiments, both implementations result in satisfactory results and can improve upon learning strategies tending to select only one generative model. Here is a summary of this work’s contributions:
  • Studying the selection task for mixtures of multiple generative models to improve the evaluation scores of generated samples (Section 4).
  • Proposing an online learning multi-armed bandit framework to address the mixture selection task for quadratic score functions (Section 5).
  • Developing the Mixture-UCB-CAB and Mixture-UCB-OGD algorithms to solve the formulated online learning problem and proving a regret bound for Mixture-UCB-CAB (Sections 5.1, 5.2).
  • Presenting numerical results on the improvements in the diversity of generated data by the online selection of a mixture of the generation models (Section 6, Appendix 8.4).
2 Related Work
Assessment of Generative Models. The evaluation of generative models has been extensively studied, with a focus on both diversity and quality of generated images. Reference-free metrics such as Rényi Kernel Entropy (RKE) (Jalali et al., 2023) and VENDI (Dan Friedman & Dieng, 2023; Ospanov et al., 2024) measure diversity without relying on ground-truth, while reference-based metrics such as Recall (Sajjadi et al., 2018; Kynkäänniemi et al., 2019) and Coverage (Naeem et al., 2020) assess diversity relative to real data. For image quality evaluation, Density and Precision metrics (Naeem et al., 2020; Kynkäänniemi et al., 2019) provide measures based on alignment with a reference distribution. The Wasserstein distance (Arjovsky et al., 2017) and Fréchet Inception Distance (FID) (Heusel et al., 2017) approximate the distance between real and generated datasets, while Kernel Inception Distance (KID) (Bińkowski et al., 2018) uses squared maximum mean discrepancy for a kernel-based comparison of distributions. Wang et al. (2023) apply the KID score for the distributed evaluation of generative models. The novelty evaluation of generative models has also been studied in References (Han et al., 2022; Jiralerspong et al., 2023; Zhang et al., 2024b; a).
Multi-Armed Bandit Algorithms. The Multi-Armed Bandit (MAB) problem is a foundational topic in reinforcement learning, where an agent aims to maximize rewards from multiple options (arms) with initially unknown reward distributions (Lai & Robbins, 1985; Thompson, 1933). The Upper Confidence Bound (UCB) algorithm (Agrawal, 1995a; Auer, 2002; Bubeck et al., 2012) is a widely adopted method for addressing the MAB problem, where uncertainty about an arm’s reward is replaced by an optimistic estimate. In generative models, optimism-based bandits have been applied to efficiently identify models with optimal Fréchet Inception Distance (FID) or Inception Score while minimizing data queries (Hu et al., 2024). A special case of MAB, the continuum-armed bandit (CAB) problem (Agrawal, 1995b), optimizes a function over continuous inputs, and has been applied to machine learning tasks such as hyperparameter optimization (Feurer & Hutter, 2019; Li et al., 2018). Recent research explores CABs under more general smoothness conditions like Besov spaces (Singh, 2021), while other works have focused on regret bounds and Lipschitz conditions (Kleinberg, 2004; Kleinberg et al., 2019; Bubeck et al., 2008).
Another related reference is informational multi-armed bandits (Weinberger & Yemini, 2023), which extends UCB to maximizing the Shannon entropy of a discrete distribution. In comparison, the algorithms in this paper can minimize the expectation of any quadratic positive-semidefinite function, which also covers the order-$2$ Rényi entropy for discrete distributions. Since the generative models’ outputs are generally continuous, (Weinberger & Yemini, 2023) is not applicable to our setting.
3 Preliminaries
We review several kernel-based performance metrics of generative models.
3.1 Rényi Kernel Entropy
The Rényi Kernel Entropy (Jalali et al., 2023) of the distribution $P$, which measures the diversity of the modes in $P$, is given by $\log(1/\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[k^{2}(X,X^{\prime})])$, where $k$ is a positive definite kernel. [^4] The order-$2$ Rényi entropy for discrete distributions is a special case by taking $k(x,x^{\prime})=\mathbf{1}_{x=x^{\prime}}$. Taking the exponential of the Rényi Kernel Entropy, we have the RKE mode count $1/\mathbb{E}[k^{2}(X,X^{\prime})])$ (Jalali et al., 2023), which is an estimate of the number of modes. Maximizing the RKE mode count is equivalent to minimizing the following loss
$\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[k^{2}(X,X^{\prime})].$ (1)
3.2 Maximum Mean Discrepancy and Kernel Inception Distance
The (squared) maximum mean discrepancy (MMD) (Gretton et al., 2012) between distributions $P,Q$, which measures the distance between $P$ and $Q$, can be written as
$\mathbb{E}[k(X,X^{\prime})]+\mathbb{E}[k(Y,Y^{\prime})]-2\mathbb{E}[k(X,Y)],$ (2)
where $X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P$ and $Y,Y^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}Q$, and $k$ is a positive definite kernel. Suppose $P$ is the distribution of samples from a generative model, and $Q$ is a reference distribution. Minimizing the MMD can ensure that $P$ is close to $Q$. The Kernel Inception Distance (KID) (Bińkowski et al., 2018), a popular quality metric for image generative models, is obtained by first passing $P$ and $Q$ through the Inception network (Szegedy et al., 2016), and then computing their MMD, i.e., we have
$\mathbb{E}[k(\psi(X),\psi(X^{\prime}))]+\mathbb{E}[k(\psi(Y),\psi(Y^{\prime}))]-2\mathbb{E}[k(\psi(X),\psi(Y))],$ (3)
where $\psi$ is the mapping from $x$ to its Inception representation.
4 Optimal Mixtures of Generative Models
RKE (1), MMD (2) and KID (3) can all be written as a loss function in the following form
$L(P):=\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[\kappa(X,X^{\prime})]+\mathbb{E}_{X\sim P}[f(X)],$ (4)
where $\kappa:\mathcal{X}^{2}\to\mathbb{R}$ is a positive semidefinite kernel, and $f:\mathcal{X}\to\mathbb{R}$ is a function. For (1), we take $\kappa(x,x^{\prime})=k^{2}(x,x^{\prime})$ (the square of a kernel is still a kernel) and $f(x)=0$. For (2), we take $\kappa(x,x^{\prime})=k(x,x^{\prime})$ and $f(x)=-2\mathbb{E}_{Y\sim Q}[k(x,Y)]$ (the constant term $\mathbb{E}[k(Y,Y^{\prime})]$ does not matter). For KID (3), we take $\kappa(x,x^{\prime})=k(\psi(x),\psi(x^{\prime}))$ and $f(x)=-2\mathbb{E}_{Y\sim Q}[k(\psi(x),\psi(Y))]$. Note that any convex combinations of (1), (2) and (3) is still in the form (4).
Suppose we are given $m$ generative models, where model $i$ generates samples from the distribution $P_{i}$. If our goal is merely to find the model that minimize the loss (4), we should select $\mathrm{argmin}_{i}L(P_{i})$. Nevertheless, for diversity metrics such as RKE, it is possible that a mixture of the models will give a better diversity. Assume that the mixture weight of model $i$ is $\alpha_{i}\in[0,1]$, where $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{m})$ is a probability vector. The loss of the mixture distribution $\sum_{i=1}^{m}\alpha_{i}P_{i}$ can then be expressed as
$L(\bm{\alpha})$ $:=L\Big{(}\sum_{i=1}^{m}\alpha_{i}P_{i}\Big{)}=\bm{\alpha}^{\top}\mathbf{K}\bm{\alpha}+\mathbf{f}^{\top}\bm{\alpha},$
$\mathbf{K}:=(\mathbb{E}_{X\sim P_{i},\,X^{\prime}\sim P_{j}}[\kappa(X,X^{\prime})])_{i,j\in[m]}\in\mathbb{R}^{m\times m},\quad\mathbf{f}:=(\mathbb{E}_{X\sim P_{i}}[f(X)])_{i=1}^{m}\in\mathbb{R}^{m}.$
Given $\mathbf{K},\mathbf{f}$, the probability vector $\bm{\alpha}$ minimizing $L(\bm{\alpha})$ can be found via a convex quadratic program.
In practice, we do not know the precise $\mathbf{K},\mathbf{f}$, and have to estimate them using samples. Suppose we have the samples $x_{i,1},\ldots,x_{i,n_{i}}$ from the distribution $P_{i}$ for $i=1,\ldots,m$, where $n_{i}$ is the number of observed samples from model $i$. Write $\mathbf{x}:=(x_{i,a})_{i\in[m],\,a\in[n_{i}]}$. We approximate the true mixture distribution $\sum_{i=1}^{m}\alpha_{i}P_{i}$ by the empirical mixture distribution $\sum_{i=1}^{m}\frac{\alpha_{i}}{n_{i}}\sum_{a=1}^{n_{i}}\delta_{x_{i,a}}$, where we assign a weight $\alpha_{i}/n_{i}$ to samples $x_{i,a}$ from model $i$, and $\delta_{x_{i,a}}$ denotes the degenerate distribution at $x_{i,a}$. We then approximate $L(\bm{\alpha})$ by the sample loss
$\hat{L}(\bm{\alpha};\mathbf{x})$ $:=L\Big{(}\sum_{i=1}^{m}\frac{\alpha_{i}}{n_{i}}\sum_{a=1}^{n_{i}}\delta_{x_{i,a}}\Big{)}=\bm{\alpha}^{\top}\hat{\mathbf{K}}(\mathbf{x})\bm{\alpha}+\hat{\mathbf{f}}(\mathbf{x})^{\top}\bm{\alpha},$ (5)
$\hat{\mathbf{K}}(\mathbf{x}):=\Big{(}\frac{1}{n_{i}n_{j}}\sum_{a=1}^{n_{i}}\sum_{b=1}^{n_{j}}\kappa(x_{i,a},x_{j,b})\Big{)}_{i,j}\in\mathbb{R}^{m\times m},\quad\hat{\mathbf{f}}(\mathbf{x}):=\Big{(}\frac{1}{n_{i}}\sum_{a=1}^{n_{i}}f(x_{i,a})\Big{)}_{i=1}^{m}\in\mathbb{R}^{m}.$
The minimization of $\hat{L}(\bm{\alpha};\mathbf{x})$ over probability vectors $\bm{\alpha}$ is still a convex quadratic program.
5 Online Selection of Optimal Mixtures – Mixture Multi-Armed Bandit
Suppose we are given $m$ generative models, but we do not have any prior information about them. Our goal is to use these models to generate a collection of samples $(x^{(t)})_{i\in[T]}$ in $T$ rounds that minimizes the loss (4) $L(\hat{P}^{(T)})$ at the empirical distribution $\hat{P}^{(T)}=T^{-1}\sum_{t=1}^{T}\delta_{x^{(t)}}$. We have
$L(\hat{P}^{(T)})=\frac{1}{T^{2}}\sum_{s,t\in[T]}\kappa(x^{(s)},x^{(t)})+\frac{1}{T}\sum_{t=1}^{T}f(x^{(t)}).$
If we are told by an oracle the optimal mixture $\bm{\alpha}^{*}$ that minimizes the loss $L(\bm{\alpha})$, then we should generate samples according to this mixture distribution, giving $\approx\alpha^{*}_{i}T$ samples from model $i$. We call this the mixture oracle scenario. Nevertheless, in reality, we do not know $\mathbf{K},\mathbf{f}$, and cannot compute $\bm{\alpha}^{*}$ exactly. Instead, we have to approximate $\bm{\alpha}^{*}$ by minimizing the sample loss $\hat{L}(\bm{\alpha};\mathbf{x})$ (5). However, we do not have the samples $\mathbf{x}$ at the beginning in order to compute $\hat{L}(\bm{\alpha};\mathbf{x})$, so we have to generate some samples first. Yet, to generate these initial samples, we need an estimate of $\bm{\alpha}^{*}$, or else those samples may have a suboptimal empirical distribution and affect our final loss $L(\hat{P}^{(T)})$, or we will have to discard those initial samples which results in wastage.
This “chicken and egg” problem is naturally solved by an online learning approach via multi-armed bandit. At time $t=1,\ldots,T$, we choose and pull an arm $b^{(t)}\in[m]$ (i.e., generate a sample from model $b^{(t)}$), and obtain a sample $x^{(t)}$ from the distribution $P_{b^{(t)}}$. The choice $b^{(t)}$ can depend on all previous samples $x^{(1)},\ldots,x^{(t-1)}$. Unlike conventional multi-armed bandit where the goal is to maximize the total reward over $T$ rounds, here we minimize the loss $L(\hat{P}^{(T)})$ which involve cross terms $\kappa(x^{(s)},x^{(t)})$ between samples at different rounds. Note that if $\kappa(x,x^{\prime})=0$, then this reduces to the conventional multi-armed bandit setting by taking $f(x)$ to be the negative reward of the sample $x$. In the following subsections, we will propose two new algorithms that are generalizations of the upper confidence bound (UCB) algorithm for multi-armed bandit (Agrawal, 1995a; Auer, 2002).
5.1 Mixture Upper Confidence Bound – Continuum-Armed Bandit
Let $n_{i}^{(t)}$ be the number of times arm $i$ has been pulled up to time $t$. Let $\mathbf{x}^{(t)}:=(x_{i,a})_{i\in[m],\,a\in[n_{i}^{(t)}]}$ be the observed samples up to time $t$. We focus on bounded loss functions, and assume that $\kappa:\mathcal{X}^{2}\to[\kappa_{0},\kappa_{1}]$ and $f:\mathcal{X}\to[f_{0},f_{1}]$ are bounded. Let $\Delta_{\kappa}:=\kappa_{1}-\kappa_{0}$, $\Delta_{f}:=f_{1}-f_{0}$. Define the sensitivity of $L$ as $\Delta_{L}:=2\Delta_{\kappa}+\Delta_{f}$.
We now present the mixture upper confidence bound – continuum-armed bandit (Mixture-UCB-CAB) algorithm. It has a parameter $\beta>1$. It treats the online selection problem as multi-armed bandit with infinitely many arms similar to the continuum-armed bandit settings in (Agrawal, 1995b; Lu et al., 2019). Each arm is a probability vector $\bm{\alpha}$. By pulling the arm $\bm{\alpha}$, we generate a sample from a randomly chosen model, where model $i$ is chosen with probability $\alpha_{i}$. Refer to Algorithm 1.
Algorithm 1 Mixture-UCB-CAB
1:  Input: $m$ generative arms, number of rounds $T$
2:  Output: Gathered samples $\mathbf{x}^{(T)}$
3:  for $t\in\{0,\ldots,m-1\}$  do
4:     Pull arm $t+1$ at time $t+1$ to obtain sample $x_{t+1,1}\sim P_{t+1}$. Set $n_{t+1}^{(m)}=1$.
5:  end for
6:  for $t\in\{m,\ldots,T-1\}$  do
7:     Compute an estimate of the optimal mixture distribution via the convex quadratic program:
$\bm{\alpha}^{(t)}:=\mathrm{argmin}_{\bm{\alpha}}\big{(}\hat{L}(\bm{\alpha};\mathbf{x}^{(t)})-(\bm{\epsilon}^{(t)})^{\top}\bm{\alpha}\big{)},$ (6)
where the minimization is over probability vectors $\bm{\alpha}$, and $\bm{\epsilon}^{(t)}\in\mathbb{R}^{m}$ is defined as
$\epsilon_{i}^{(t)}:=\Delta_{L}\sqrt{(\beta\log t)/(2n_{i}^{(t)})}+\Delta_{\kappa}/n_{i}^{(t)}.$ (7)
8:     Generate the arm index $b^{(t+1)}\in[m]$ at random with $\mathbb{P}(b^{(t+1)}=i)=\alpha_{i}^{(t)}$.
9:     Pull arm $b=b^{(t+1)}$ at time $t+1$ to obtain a new sample $x_{b,n_{b}^{(t)}+1}\sim P_{b}$. Set $n_{b}^{(t+1)}=n_{b}^{(t)}+1$ and $n_{j}^{(t+1)}=n_{j}^{(t)}$ for $j\neq b$55footnotemark: 5
10:  end for
11:  return  samples $\mathbf{x}^{(T)}$
Similar to UCB, Mixture-UCB-CAB finds a lower confidence bound $\hat{L}(\bm{\alpha};\mathbf{x}^{(t)})-(\bm{\epsilon}^{(t)})^{\top}\bm{\alpha}$ of the true loss $L(\bm{\alpha})$ at each round. To justify the expressions (6), (7), we prove that $\hat{L}(\bm{\alpha};\mathbf{x}^{(t)})-(\bm{\epsilon}^{(t)})^{\top}\bm{\alpha}$ in (6) lower-bounds $L(\bm{\alpha})$ with probability at least $1-t^{-\beta}$. The proof is given in Appendix 8.1.
Theorem 1
Fix a probability vector $\bm{\alpha}$ [^6] Theorem 1 holds for a fixed $\bm{\alpha}$. A worst-case bound that simultaneously holds for every $\bm{\alpha}$ is in Lemma 1. Suppose we have samples $x_{i,1},\ldots,x_{i,n_{i}}$ from the distribution $P_{i}$ for $i=1,\ldots,m$, where $n_{i}$ is the number of observed samples from model $i$. For $\delta>0$,
$\mathbb{P}\big{(}\hat{L}(\bm{\alpha};\mathbf{x})-L(\bm{\alpha})\geq\bm{\epsilon}(\delta)^{\top}\bm{\alpha}\big{)}\leq\delta,\quad\mathbb{P}\big{(}L(\bm{\alpha})-\hat{L}(\bm{\alpha};\mathbf{x})\geq\bm{\epsilon}(\delta)^{\top}\bm{\alpha}\big{)}\leq\delta,$
where $\bm{\epsilon}(\delta):=\big{(}\Delta_{L}\sqrt{\frac{\log(1/\delta)}{2n_{i}}}+\frac{\Delta_{\kappa}}{n_{i}}\big{)}_{i\in[m]}$.
We now prove that Mixture-UCB-CAB gives an expected loss $\mathbb{E}[L(\hat{P}^{(T)})]$ that converges to the optimal loss $\min_{\bm{\alpha}}L(\bm{\alpha})$ by bounding their gap. This means that Mixture-UCB-CAB is a zero-regret strategy by treating $\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\bm{\alpha}}L(\bm{\alpha})$ as the average regret per round. [^7] To justify calling $R:=\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\bm{\alpha}}L(\bm{\alpha})$ the average regret per round, note that when $\kappa(x,x^{\prime})=0$ and $f(x)=-r(x)$ where $r(x)$ is the reward of the sample $x$, i.e., the loss $L(P)=\mathbb{E}_{X\sim P}[f(X)]$ is linear, $T(\mathbb{E}[L(\hat{P}^{(T)})]-\min_{\bm{\alpha}}L(\bm{\alpha}))=T\max_{i\in[m]}\mathbb{E}_{X\sim P_{i}}[r(X)]-\mathbb{E}[\sum_{t=1}^{T}r(x^{(t)})]$ indeed reduces to the conventional notion of regret. So $R$ can be regarded as the quadratic generalization of regret. The proof is given in Appendix 8.2.
Theorem 2
Suppose $m\geq 2$, $\beta\geq 4$. Consider bounded quadratic loss function (4) with $\kappa$ being positive semidefinite. Let $\hat{P}^{(T)}$ be the empirical distribution of the first $T\geq 2$ samples $\mathbf{x}^{(T)}$ given by Mixture-UCB-CAB. Then the gap between the expected loss and the optimal loss is bounded by
$\mathbb{E}\big{[}L(\hat{P}^{(T)})\big{]}-\min_{\bm{\alpha}}L(\bm{\alpha})$ $\leq 4\Delta_{L}\sqrt{\frac{\beta m\log T}{T}}.$
When $\kappa(x,x^{\prime})=0$, Mixture-UCB-CAB reduces to the conventional UCB, and Theorem 2 coincides with the $O(\sqrt{(m\log T)/T})$ distribution-free bound on the regret per round of conventional UCB (Bubeck et al., 2012). Since there is a $\Omega(\sqrt{m/T})$ minimax lower bound on the regret per round even for conventional multi-armed bandit without the quadratic kernel term (Bubeck et al., 2012, Theorem 3.4), Theorem 2 is tight up to a logarithmic factor.
The main difference between Mixture-UCB-CAB and conventional UCB is that we choose a mixture of arms in (6) given by the probability vector $\bm{\alpha}$, instead of a single arm. A more straightforward application of UCB would be to simply find the single arm that minimizes the lower bound in (6), i.e., we restrict $\bm{\alpha}=\mathbf{e}_{i}$ for some $i\in[m]$, where $\mathbf{e}_{i}$ is the $i$-th basis vector, and minimize (6) over $i$ instead. We call this Vanilla-UCB. Vanilla-UCB fails to take into account the possibility that a mixture may give a smaller loss than every single arm. In the long run, Vanilla-UCB converges to pulling the best single arm instead of the optimal mixture. Vanilla-UCB will be used as a baseline to be compared with Mixture-UCB-CAB, and another new algorithm presented in the next section.
Mixture-UCB-CAB can be extended to the Sparse-Mixture-UCB-CAB algorithm which eventually select only a small number of models. This can be useful if there is a subscription cost for each model. Refer to Appendix 8.3 for discussions.
77footnotetext: We may also consider the scenario where each pull gives a batch of $l$ samples instead of only one sample. In this case, we will have $x_{b,n_{b}^{(t-1)}+1},\ldots,x_{b,n_{b}^{(t-1)}+l}\sim P_{b}$ and $n_{b}^{(t)}=n_{b}^{(t-1)}+l$.
5.2 Mixture Upper Confidence Bound – Online Gradient Descent
We present an alternative to Mixture-UCB-CAB, called the mixture upper confidence bound – online gradient descent (Mixture-UCB-OGD) algorithm, inspired by the online gradient descent algorithm (Shalev-Shwartz et al., 2012). It also has a parameter $\beta>1$. Refer to Algorithm 2.
Algorithm 2 Mixture-UCB-OGD
1:  Input: $m$ generative arms, number of rounds $T$
2:  Output: Gathered samples $\mathbf{x}^{(T)}$
3:  for $t\in\{0,\ldots,m-1\}$  do
4:     Pull arm $t+1$ at time $t+1$ to obtain sample $x_{t+1,1}\sim P_{t+1}$
5:  end for
6:  for $t\in\{m,\ldots,T-1\}$  do
7:     Compute the gradient
$\mathbf{h}^{(t)}$ $:=\nabla_{\bm{\alpha}}\left(\hat{L}(\bm{\alpha};\mathbf{x}^{(t)})-(\bm{\epsilon}^{(t)})^{\top}\bm{\alpha}\right)\Big{|}_{\bm{\alpha}=\mathbf{n}^{(t)}/t}\;=\;\frac{2}{t}\hat{\mathbf{K}}(\mathbf{x}^{(t)})\mathbf{n}^{(t)}+\hat{\mathbf{f}}(\mathbf{x}^{(t)})-\bm{\epsilon}^{(t)},$ (8)
where $\mathbf{n}^{(t)}:=(n_{i}^{(t)})_{i\in[m]}\in\mathbb{R}^{m}$, and $\bm{\epsilon}^{(t)}$ is defined as in Mixture-UCB-CAB
8:     Pull arm $b=b^{(t+1)}:=\mathrm{argmin}_{i}h_{i}^{(t)}$ at time $t+1$ to obtain a new sample $x_{b,n_{b}^{(t)}+1}\sim P_{b}$.
9:  end for
10:  return  samples $\mathbf{x}^{(T)}$
Mixture-UCB-CAB and Mixture-UCB-OGD can both be regarded as generalizations of the original UCB algorithm, in the sense that they reduce to UCB when $\kappa(x,x^{\prime})=0$. If we remove the $\frac{2}{t}\hat{\mathbf{K}}(\mathbf{x})\mathbf{n}^{(t)}$ term in (8), then Mixture-UCB-OGD becomes the same as UCB.
Both Mixture-UCB-CAB and Mixture-UCB-OGD attempt to make the “proportion vector” $\mathbf{n}^{(t)}/t$ (note that $n_{i}^{(t)}/t$ is the proportion of samples from model $i$) approach the optimal mixture $\bm{\alpha}^{*}$ that minimizes $L(\bm{\alpha})$, but they do so in different manners. Mixture-UCB-CAB first computes the estimate $\bm{\alpha}^{(t)}$ after time $t$, then approaches $\bm{\alpha}^{(t)}$ by pulling an arm randomly chosen from the distribution $\bm{\alpha}^{(t)}$. Mixture-UCB-OGD estimates the gradient $\mathbf{h}^{(t)}$ of the loss function at the current proportion vector $\mathbf{n}^{(t)}/t$, and pulls an arm that results in the steepest descent of the loss.
An advantage of Mixture-UCB-OGD is that the computation of gradient (8) is significantly faster than the quadratic program (6) in Mixture-UCB-CAB. The running time complexity of Mixture-UCB-OGD is $O(T^{2}+Tm^{2})$. [^8] To update $\hat{\mathbf{K}}(\mathbf{x}^{(t)})$ after a new sample $x^{\prime}$ is obtained, we only need to compute $\kappa(x,x^{\prime})$ for each existing sample $x$, and add their contributions to the corresponding entries in $\hat{\mathbf{K}}(\mathbf{x}^{(t)})$, requiring a computational time that is linear with the number of existing samples. Nevertheless, a regret bound for Mixture-UCB-OGD similar to Theorem 2 seems to be difficult to derive, and is left for future research.
6 Numerical Results
We experiment on various scenarios to showcase the performance of our proposed algorithms. The experiments involve the following algorithms:
  • Mixture Oracle. In the mixture oracle algorithm (Section 5), an oracle tells us the optimal mixture $\bm{\alpha}^{*}$ in advance, and we pull arms randomly according to this distribution. The optimal mixture is calculated by solving the quadratic optimization in Section 4 on a large number of samples for each arm. The number of chosen samples varies based on the experiments. This is an unrealistic setting that only serves as a theoretical upper bound of the performance of any online algorithm. A realistic algorithm that performs close to the mixture oracle would be almost optimal.
  • One-Arm Oracle. An oracle tells us the optimal single arm in advance, and we keep pulling this arm. This is an unrealistic setting. If our algorithms outperform the one-arm oracle, this will show that the advantage of pulling a mixture of arms (instead of a single arm) can be realistically achieved via online algorithms.
  • Vanilla-UCB. A direct application of UCB mentioned near the end of Section 5.1. This serves as a baseline for the purpose of comparison.
  • Successive Halving. The Success Halving algorithm (Karnin et al., 2013; Jamieson & Talwalkar, 2016; Chen & Ghosh, 2024) which serves as a second baseline for comparison.
  • Mixture-UCB-CAB. The mixture upper confidence bound – continuum-armed bandit algorithm proposed in Section 5.1.
  • Mixture-UCB-OGD. The mixture upper confidence bound – online gradient descent algorithm proposed in Section 5.2.
Experiments Setup.
We used DINOv2-ViT-L/14 (Oquab et al., 2023) for image feature extraction, as recommended in (Stein et al., 2023), and RoBERTa (Liu et al., 2019) as the text encoder. Detailed explanation of the setup for each experiment is presented in Section 8.4.
6.1 Optimal Mixture for Diversity and Quality via KID
(a) FFHQ generators
(b) LSUN-Bedroom generators
(c) FFHQ truncated generators
Figure 3: Performance comparison of online algorithms for the KID metric across FFHQ, LSUN-Bedroom, and FFHQ Truncated generators.
We conducted three experiments to evaluate our method using the Kernel Inception Distance (KID) metric. In the first experiment, we used five distinct generative models: LDM (Rombach et al., 2022), StyleGAN-XL (Sauer et al., 2022), Efficient-VDVAE (Hazami et al., 2022), InsGen (Yang et al., 2021), and StyleNAT (Walton et al., 2022), all trained on the FFHQ dataset (Karras et al., 2019). In the second experiment, we used generated images from four models [^9] FFHQ and LSUN-Bedroom datasets were downloaded from the dgm-eval repository (Stein et al., 2023) (licensed under MIT license): https://github.com/layer6ai-labs/dgm-eval.: StyleGAN (Karras et al., 2019), Projected GAN (Sauer et al., 2021), iDDPM (Nichol & Dhariwal, 2021), and Unleashing Transformers (Bond-Taylor et al., 2022), all trained on the LSUN-Bedroom dataset (Yu et al., 2015). This experiment followed a similar setup to the first. In the final experiment, we employed the truncation method (Marchesi, 2017; Karras et al., 2019) to generate diversity-controlled images centered on eight randomly selected points, using StyleGAN2-ADA (Karras et al., 2020), also trained on the FFHQ dataset. Figure 3 demonstrates that the mixture of generators achieves better KID scores compared to individual models. Additionally, the two Mixture-UCB algorithms consistently outperform the baselines.
6.2 Optimal Mixture for Diversity via RKE
We used the RKE Mode Count (Jalali et al., 2023) as an evaluation metric to show the effect of mixing the models on the diversity and the advantage of our algorithms Mixture-UCB-CAB and Mixture-UCB-OGD. The score in the plots is the RKE Mode Count, written as RKE for brevity.
Synthetic Unconditional Generative Models
We conduct two experiments on diversity-limited generative models. First, we used eight center points with a truncation value of 0.3 to generate images using StyleGAN2-ADA, trained on the FFHQ dataset. In the second experiment, we applied the same model, trained on the AFHQ Cat dataset (Choi et al., 2020), with a truncation value of 0.4. As shown in Figure 7, the optimal mixture and our algorithms consistently achieve higher RKE scores. The increase in diversity is visually depicted in Figures 6 and 8.
Text to Image Generative Models
We used Stable Diffusion XL (Podell et al., 2024) with specific prompts to create three car image generators with distinct styles: realistic, surreal, and cartoon. In the second experiment, recognizing the importance of diversity in generative models for design tasks, we used five models—FLUX.1-Schnell (Labs, 2024), Kandinsky 3.0 (Vladimir et al., 2024), PixArt-$\alpha$ (Chen et al., 2023a), and Stable Diffusion XL—to generate images of the object “Sofa”. In a similar manner, we generated red bird images using Kandinsky 3.0, Stable Diffusion 3 (Esser et al., 2024), and PixArt-$\alpha$, as shown in Figure 2.. Finally, in the third experiment, we used Stable Diffusion XL to simulate models generating images of different dog breeds: Bulldog, German Shepherd, and Poodle, respectively. This illustrates the challenge of generating diverse object types with text-to-image models. Figure 10 demonstrates the impact of using a mixture of models in the first and third experiments. The improvement in diversity is evident visually and quantitatively, as shown by the RKE scores. Our online algorithms consistently generate more diverse samples than others, as illustrated in Figure 4.
(a) Dog breed generators
(b) Sofa image generators
(c) Style-specific generators
Figure 4: Performance comparison of online algorithms using RKE score of T2I generative models.
6.3 Optimal Mixture for Diversity and Quality via RKE and Precision/Density
Using RKE, we focus solely on the diversity of the arms without accounting for their quality. To address this, we apply our methodology to both RKE and Precision (Kynkäänniemi et al., 2019), as well as RKE and Density (Naeem et al., 2020). We conduct experiments in which quality is a key consideration. We use four arms: three are StyleGAN2-ADA models trained on the FFHQ dataset, each generating images with a truncation of 0.3 around randomly selected center points. The fourth model is StyleGAN2-ADA trained on CIFAR-10 (Krizhevsky et al., 2009). The FFHQ dataset is used as the reference dataset. Figures 5 and 13 demonstrate the ability of our algorithms in finding optimal mixtures with higher diversity/quality score.
(a) RKE w/Precision
(b) RKE w/Density
Figure 5: Performance comparison of online algorithms using the combination of RKE with Precision and RKE with Density metrics.
7 Conclusion and Limitations
We studied the online selection from several generative models, where the online learner aims to generate samples with the best overall quality and diversity. While standard multi-armed bandit (MAB) algorithms converge to one arm and select one generative model, we highlighted that a mixture of generative models could achieve a higher score compared to the individual models. We proposed the Mixture-UCB MAB algorithm to find the optimal mixture. Our experiments suggest the usefulness of the algorithm in improving the performance scores over individual arms. However, we note that the diversity gain offered by the mixture approach should be analyzed together with the quality of generated data, which can be adjusted by setting the coefficient of the quality Precision or Density score when applying Mixture-UCB. The analysis of the regret of Mixture-UCB-OGD and conditions under which a mixture can or cannot improve the scores will remain for future studies.
Acknowledgments
The work of Cheuk Ting Li is partially supported by two grants from the Research Grants Council of the Hong Kong Special Administrative Region, China [Project No.s: CUHK 24205621 (ECS), CUHK 14209823 (GRF)]. The work of Farzan Farnia is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by CUHK Direct Research Grants with CUHK Project No. 4055164 and 4937054. Finally, the authors would like to thank the anonymous reviewers for their constructive feedback and suggestions.